### v2

# Provable Emergence of Deep Neural Collapse and Low-Rank Bias in $L^{2}$-Regularized Nonlinear Networks

Abstract
Recent work in deep learning has shown strong empirical and theoretical evidence of an implicit low-rank bias: weight matrices in deep networks tend to be approximately low-rank. Moreover, removing relatively small singular values during training, or from available trained models, may significantly reduce model size while maintaining or even improving model performance. However, the majority of the theoretical investigations around low-rank bias in neural networks deal with oversimplified models, often not taking into account the impact of nonlinearity. In this work, we first of all quantify a link between the phenomenon of deep neural collapse and the emergence of low-rank weight matrices for a general class of feedforward networks with nonlinear activation. In addition, for the general class of nonlinear feedforward and residual networks, we prove the global optimality of deep neural collapsed configurations and the practical absence of a loss barrier between interpolating minima and globally optimal points, offering a possible explanation for its common occurrence. As a byproduct, our theory also allows us to forecast the final global structure of singular values before training. Our theoretical findings are supported by a range of experimental evaluations illustrating the phenomenon.
1 Introduction and related work
Recent years have seen growing theoretical and experimental evidence of low-rank properties of deep neural networks. For example, it has been noted that a large portion of singular values from linear layers in different models can be removed almost without affecting performance, both for feedforward and residual networks Schotthofer_low_rank ; Ydelbayev_low_rank . Analogous results have been shown on feedforward and convolutional networks for image classification Li_basis_CNN ; wang2021pufferfish ; khodak2021initialization ; Schotthofer_low_rank as well as on Recurrent Neural Networks (RNNs) trained to learn neural dynamics in neuroscience tasks schuessler2020interplay ; pellegrino2023low .
Alongside the wealth of experimental evidence of low-rank properties of neural networks, several recent works have investigated the low-rank bias phenomenon from the theoretical point of view, providing evidence of decaying singular values in deep linear architectures arora2019implicit ; feng2022rank ; huh2022low and of low-rank properties of networks’ Hessian maps singh2021analytic , analyzing low-rank geometric properties of training gradient flow dynamics for deep linear networks bah2022learning , showing that low-rank bias is related to weight decay galanti2023characterizing ; sukenik2024neural ; kuzborskij2025lowrankbiasweightdecay , and to training invariances le2022training ; ziyin2025parametersymmetrybreakingrestoration . In particular, the authors relate the rank of the weight matrices during SGD to the batch size adopted during the training of the network. However, based only on the batch size, the results from galanti2023characterizing become uninformative when the batch size is comparable to the number of parameters of the neural network.
Together with this, another related phenomenon often observed during the training of classifiers is the so-called “neural collapse” papyan2020prevalence ; zhu2021geometric , which is often described using four properties named NC1–4. Different works have been proving the optimality of neural collapse in different simplified settings zhu2021geometric ; seleznova2023neural ; jacot2025wide , and different extensions have been proposed for its deep version. In particular, we are interested in the first property of Deep Neural Collapse ($\operatorname{\mathrm{DNC1}}$), which essentially states that the within-class variability of intermediate representations goes to zero even for layers in the middle of a deep neural network rangamani2023feature ; sukenik2024neural ; sukenik2024deep (see Definition˜3.5).
In this work, we show a connection between the phenomenon of deep neural collapse and low-rank implicit bias in general nonlinear networks. The main contribution is theoretical and in particular, our main results, which hold in the presence of weight decay, can be summarized as follows:
  • in Section˜3 we prove that within-class variability of hidden layers during training (including the initial dataset) has a quantifiable effect on the numerical rank of the network parameters as the latter decays with the total cluster variation of the hidden representations;
  • in Section˜4 we show that zero total cluster variation for all intermediate layers is optimal in a fairly general setting, both for feedforward and residual networks (Theorem˜4.2). Moreover, in Theorem˜4.3, we prove that for (almost) every interpolating initial condition, there exists a loss-decreasing trajectory connecting the initial condition to a collapsed configuration, thereby offering an explanation to the consistent empirical observation of deep neural collapse rangamani2023feature .
Overall, the combination of the results above proves that in the presence of a small weight decay regularization, zero within-class variability for all intermediate representations is globally optimal under some natural constraints, therefore showing that global minima in parameter space are low-rank matrices. Moreover, convergence to these highly structured solutions may be regarded as universal in the sense that, in the presence of weight decay there is essentially no loss barrier between interpolating features to deep-neural collapsed ones.
The key points of our investigation are the following observations:
  • The stationary points of weight decay-regularized losses $\mathcal{L}_{\lambda}(\mathit{\Theta},x)=\mathcal{L}_{0}(\mathit{\Theta},x)+\lambda\|\mathit{\Theta}\|^{2}$ are solutions to nonlinear eigenvalue equations
    $\nabla_{W_{l}}\mathcal{L}_{0}(\mathit{\Theta},x)=-\lambda W_{l},$
    where $W_{l}$ is the weight matrix of the $l$-th layer. This eigenvalue equation allows us to transfer properties from the gradient to the parameters and vice versa. A similar connection between parameter ranks and weight decay has recently been used in galanti2023characterizing to address the implicit bias problem from a different perspective.
  • The rank of the gradient of the nonregularized loss $\mathcal{L}$ evaluated on $k$ training examples is controlled by $k$, namely $\operatorname{rank}\nabla_{W_{l}}\mathcal{L}_{0}(\mathit{\Theta},x)\leq k$.
  • For a sequence of layers with weight-decay, it is possible to reparameterize the representation cost as a function of the hidden features, as done in jacot2022featurelearningl2regularizeddnns . This setting is not restrictive since the stationary points of the reparametrized cost are in bijection with the original ones jacot2022featurelearningl2regularizeddnns .
  • It is possible to study exactly the global minima of the representation cost $\mathscr{L}(\mathit{\Theta})=\sum_{l=2}^{L}\|W_{l}\|^{2}$ constrained to the set of interpolating configurations, proving that all intermediate features satisfy $\operatorname{\mathrm{DNC1}}$. Thanks to this last result, it is also possible to study the structure of the interpolating minima of $\sum_{l=1}^{L}\|W_{l}\|^{2}$ as a perturbative limit of the previous one when the norm of the input data is much bigger than the norm of the target matrix, i.e., $\|X\|\gg\|Y\|$ (Proposition˜A.2), or for general data matrices in the large depth regime (Remark˜B.6).
  • In this setting, it is also possible to study mode connectivity of the loss landscape under a set of natural constraints $\mathcal{S}$. In particular, we prove that for (almost) every $\mathit{\Theta}(0)$ such that $\mathcal{L}_{0}(\mathit{\Theta}(0),X)=0$ we have:
    $\inf_{\gamma\in\Gamma}\sup_{t\in[0,1]}\mathscr{L}(\gamma(t))\leq\mathscr{L}(\mathit{\Theta}(0))$
    where $\Gamma:=\{\gamma\in\mathcal{C}^{0}([0,1];\mathcal{S})\,|\,:\gamma(0)=\mathit{\Theta}(0),\gamma(1)\,\text{satisfies}\,\,\operatorname{\mathrm{DNC1}}\,\forall l=1,\dots,L\}$.
Together, these results show that there is essentially no loss barrier between zero loss points and global minima of $\mathscr{L}$, which we prove are all deep-neural collapsed. This will be proven constructively, showing the existence of an energy decreasing path with the desired property. Being this result decoupled from the particular dynamics during training, it might be a step in explaining the universality in the emergence of (deep) neural collapse. Moreover, we highlight that, despite the convexity of $\mathscr{L}$, the constraint $\mathcal{L}_{0}(\mathit{\Theta};X)=0$ is highly nonconvex, making this last result nontrivial.
2 Setting and preliminaries
Consider a dataset $\mathcal{X}=\{x_{j}\}_{j=1}^{N}\subseteq\mathbb{R}^{d}$ with $c$ corresponding labels $\mathcal{Y}=\{y_{j}\}_{j=1}^{N}\subseteq\mathbb{R}^{c}$.
We assume to train a fully connected $L$-layer feedforward neural network $f(\mathit{\Theta},x)$ of the form
$z^{0}=x,\quad f(\mathit{\Theta},x)=z^{L},\quad z^{l}=\sigma_{l}\big{(}W_{l}z^{l-1}+b_{l}\big{)},\;l=1,\dots,L$ (1)
to learn from the given data and labels. In equation (1), $\sigma_{l}$ are scalar nonlinear activation functions acting entrywise and $\mathit{\Theta}$ denotes all trainable parameters $\mathit{\Theta}=(W_{1},\dots,W_{L},b_{1},\dots,b_{L})$. We write
$\mathit{\Theta}_{l_{1}}^{l_{2}}$ $=(W_{l_{1}+1},\dots,W_{l_{2}},b_{l_{1}+1},\dots,b_{l_{2}}),\quad l_{1}<l_{2},$ (2)
$f_{l_{1}}^{l_{2}}(\mathit{\Theta}_{l_{1}}^{l_{2}},z^{l_{1}})$ $=\sigma(W_{l_{2}}\cdots\sigma(W_{l_{1}+1}z^{l_{1}}+b_{l_{1}+1})+\cdots+b_{l_{2}})$
to denote the training parameters and concatenation of layers from $l_{1}+1$ to $l_{2}$, respectively. We adopt the convention that, when we do not specify the lower index then it starts from zero, and when we omit the top one it ends at the overall depth $L$, namely, $\mathit{\Theta}^{l_{2}}:=\mathit{\Theta}_{0}^{l_{2}}$, $f^{l_{2}}:=f^{l_{2}}_{0}$ and $\mathit{\Theta}_{l_{1}}:=\mathit{\Theta}_{l_{1}}^{L}$, $f_{l_{1}}:=f^{L}_{l_{1}}$. Notice that, with this notation, $\mathit{\Theta}=\mathit{\Theta}_{0}=\mathit{\Theta}^{L}=\mathit{\Theta}_{0}^{L}$, $f=f_{0}=f^{L}=f_{0}^{L}$ and $\mathit{\Theta}^{0}$ is empty. Finally, for the sake of brevity, we write $\mathcal{Z}^{l}:=\{z^{l}_{i}\}_{x_{i}\in\mathcal{X}}$ to denote the set of outputs of the $l$-th layer, where $z^{l}_{i}=f^{l}(\mathit{\Theta}^{l},x_{i})$ is defined as in (1) with input data $x_{i}$, and $\mathcal{Z}^{0}=\mathcal{X}$.
Moreover, to simplify notation, we will denote the feature matrices of input, output, and intermediate layers as
$X=[x_{1},\dots,x_{N}]\in\mathbb{R}^{d\times N},\quad Y=[y_{1},\dots,y_{N}]\in\mathbb{R}^{c\times N},\quad Z^{l}=[z_{1}^{l},\dots,z_{N}^{l}]\in\mathbb{R}^{n_{l}\times N},$
where $1\leq l\leq L$. The parameters $\mathit{\Theta}$ are assumed to be trained by minimizing a regularized loss function $\operatorname{\mathcal{L}}_{\lambda}(\mathit{\Theta})$ composed by the sum of losses at individual data points, i.e.,
$\operatorname{\mathcal{L}}_{\lambda}(\mathit{\Theta}):=\frac{1}{N}\sum_{i=1}^{N}\operatorname{\ell}\big{(}f(\mathit{\Theta},x_{i}),y_{i}\big{)}+\lambda\|\mathit{\Theta}\|^{2},\quad\text{where}\quad\|\mathit{\Theta}\|^{2}=\sum_{i=1}^{L}\Bigl{(}\|W_{i}\|^{2}+\|b_{i}\|^{2}\Bigr{)}.$ (3)
Here $\|\cdot\|$ denotes the entrywise $L^{2}$ norm (when the input is a matrix, it coincides with the Frobenius norm) and $\lambda\geq 0$ is a weight decay or regularization parameter.
In this setting, in Section˜3 we will quantify a precise relationship between the numerical rank of the stationary points of $\mathcal{L}_{\lambda}$ and the total cluster variation of intermediate feature matrices.
3 Neural Rank Collapse: a link between low-rank and deep-neural collapse
Note that any critical point of $\mathcal{L}_{\lambda}(\mathit{\Theta})$ satisfies the following equation:
$\lambda W_{l}^{*}=-\nabla_{W_{l}}\operatorname{\mathcal{L}}_{0}(\mathit{\Theta}^{*})\qquad\forall l=1,\dots,L.$ (4)
Equation (4) corresponds to an “eigenvalue equation” for the nonlinear operator $\nabla\operatorname{\mathcal{L}}_{0}$. In other words, for a fixed weight decay parameter $\lambda$, training the neural network corresponds to finding a nonlinear eigenvector $\mathit{\Theta}^{*}$ having $\lambda$ as eigenvalue. In addition, equation (4) immediately implies that whenever the operators $\nabla_{W_{l}}\operatorname{\mathcal{L}}_{0}$ have low rank (i.e., they take values in some low-rank space), then also any weight matrix $W^{*}$ composing the eigenvector $\mathit{\Theta}^{*}$ has necessarily low rank. In the following, we investigate this relationship more in depth, providing sufficient conditions to have approximately low-rank gradients.
[…]
Proposition 3.1 (Small batches yield low-rank gradients).
Let $f(\mathit{\Theta},x)$ be a feedforward neural network as in (1) and $\operatorname{\ell}:\mathbb{R}^{c}\times\mathbb{R}^{c}\to\mathbb{R}$ be a differentiable loss function as in (3). Then, for any $\{(x_{i},y_{i})\}_{i=1}^{K}\subseteq\mathbb{R}^{d}\times\mathbb{R}^{c}$, we have $\operatorname{rank}\Big{(}\operatorname{\nabla}_{W_{l}}\sum_{i=1}^{K}\operatorname{\ell}\big{(}f(\mathit{\Theta},x_{i}),y_{i}\big{)}\Big{)}\leq K$.
Note that, given $z^{j}=f^{j}(\mathit{\Theta}^{j},x)$, as long as $j<l$, it holds
$\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f(\mathit{\Theta},x),y\big{)}=\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},z^{j}),y\big{)}.$ (5)
[…]
In addition, Proposition˜3.1 immediately yields an example where gradients have necessarily low rank. Consider the case of degenerate data, i.e., assume that there exist $K$ pairs $\{(\bar{x}_{k},\bar{y}_{k})\}_{k=1}^{K}$ such that any data point $(x_{i},y_{i})$ is equal to some $(\bar{x}_{k},\bar{y}_{k})$. Then, by Proposition˜3.1, $\operatorname{\nabla}_{W_{l}}\sum_{i}\operatorname{\ell}(\mathit{\Theta},x_{i},y_{i})$ has rank at most $K$. But the right-hand side of (4) has rank at most $K$, hence any critical point of the loss function $\operatorname{\mathcal{L}}_{\lambda}$ necessarily corresponds to a neural network having all layers of rank at most $K$. The same argument can be extended to the situation of total intermediate neural collapse Poggio_hidden_nc , where the intermediate hidden set $\mathcal{Z}^{l}$ consists of at most $K$ distinct points. Indeed, assume that for a stationary point $\mathit{\Theta}^{*}$ there exists a layer $l$ and points $\{(\bar{z}_{k}^{l},\bar{y}_{k})\}_{k=1}^{K}$ such that any pair $(z^{l}_{i},y_{i})=(\bar{z}_{k}^{l},\bar{y}_{k})$ for some $k$. Then, by Proposition˜3.1 and the subsequent remark (5), any subsequent layer, $W_{j}^{*}$ with $j>l$, has rank at most $K$.
In the following, we prove that under weaker hypotheses, like the collapse of hidden representations, any critical point of the loss function produces layers close to low-rank. In particular, the approximate rank of the layers at a stationary point is given by the number of clusters (or collapsing sets) and the approximation error depends both on the regularity of the loss function in the clusters and the within-cluster variability, which we measure as the total cluster variation of each hidden set $\mathcal{Z}^{l}$.
3.1 Clustered Data and Total Cluster Variation
Assume $\mathcal{C}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{K}\}$ to be a $K$-partition of the input data $\mathcal{X}$ and corresponding labels $\mathcal{Y}$, i.e., for any $(x_{i},y_{i})\in\mathcal{X}\times\mathcal{Y}$ there exists a unique $\mathcal{C}_{k}\in\mathcal{C}$ such that $(x_{i},y_{i})\in\mathcal{C}_{k}$. We denote by $N_{k}=|\mathcal{C}_{k}|$ the cardinality of the $k$-th family and adopt the notation
$\mathcal{X}\times\mathcal{Y}=\bigcup_{k=1}^{K}\bigcup_{i\in\mathcal{C}_{k}}\{(x_{i},y_{i})\}=\bigcup_{k=1}^{K}\bigcup_{i=1}^{N_{k}}\{(x_{i}^{k},y_{i}^{k})\},$
where with a small abuse of notation we use $\mathcal{C}_{k}$ to denote also the set of indices whose corresponding data points are contained in $\mathcal{C}_{k}$. Now let $\bar{x}_{k}=\bar{z}^{0}_{k}$, $\bar{y}_{k}$ and $\bar{z}^{l}_{k}$ be the centroids of the sets $\{x_{i}\}_{i\in\mathcal{C}_{k}}$, $\{y_{i}\}_{i\in\mathcal{C}_{k}}$ and $\{z^{l}_{i}\}_{i\in\mathcal{C}_{k}}$, respectively, i.e.,
$\bar{z}^{l}_{k}=\frac{1}{N_{K}}\sum_{i\in\mathcal{C}_{k}}z_{i}^{l},\qquad\bar{y}_{k}=\frac{1}{N_{K}}\sum_{i\in\mathcal{C}_{k}}y_{i}.$
To quantify how clustered the data is with respect to the partition $\mathcal{C}$, we consider the Within-Cluster Sum of Squares (WCSS), defined as
$\operatorname{\mathrm{WCSS}}_{l}(\mathcal{C})=\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in\mathcal{C}_{k}}\|\big{(}z^{l}_{i},y_{i}\big{)}-\big{(}\bar{z}_{k}^{l},\bar{y}_{k}\big{)}\|^{2}.$ (6)
This quantity measures the within-cluster variance (see, e.g., hastie2009elements ) of the data at layer $l$ with respect to a prescribed partition and the corresponding centroids.
Note that having a clustered dataset at the $l$-th layer means that there exists a suitable partition $\mathcal{C}^{*}$ such that $\operatorname{\mathrm{WCSS}}_{l}(\mathcal{C}^{*})$ is sufficiently small. To measure the within-cluster variability of the data at the $l$-th layer when it is partitioned into at most $K$ clusters, we introduce the Total Cluster Variation.
Definition 3.2 (Total Cluster Variation).
The $K$-th Total Cluster Variation (TCV) at layer $l$, is defined as
$\operatorname{\mathrm{TCV}}_{K,l}:=\min_{\mathcal{C}\in\cup_{r=1}^{K}\mathcal{P}_{r}}\operatorname{\mathrm{WCSS}}_{l}(\mathcal{C}),$ (7)
where $\mathcal{P}_{r}$ is the set of all partitions of $\{(z_{i}^{l},y_{i})\}_{i=1}^{N}$ into $r$ sets.
Note that, for $K$ sufficiently smaller than $N$, a small value of $\operatorname{\mathrm{TCV}}_{K,l}$ indicates an intermediate collapse of the network’s features at layer $l$. Moreover, being $\operatorname{\mathrm{TCV}}_{K,l}$ defined through a variational problem to get the tightest bound possible, it is trivially upper bounded by the total cluster variation on the classes. Therefore, neural collapse in the classical sense implies $\operatorname{\mathrm{TCV}}_{K,l}=0$.
3.2 Neural collapse yields low-rank bias
In this section, we show that the occurrence of well-clustered hidden representations, i.e., of intermediate neural collapse, necessarily yields low-rank properties for the underlying neural network in the presence of weight decay. Given a partition $\mathcal{C}=\{\mathcal{C}_{k}\}_{k=1}^{K}$ as in Section˜3.1, we can introduce the centroid-based loss at the $l$-th layer:
$\operatorname{\mathcal{L}}^{l,\mathcal{C}}_{\lambda}(\mathit{\Theta})=\sum_{k=1}^{K}\pi_{k}\operatorname{\ell}\big{(}f_{l}(\mathit{\Theta}_{l},\bar{z}_{k}^{l}),\bar{y}_{k}\big{)}+\lambda\|\mathit{\Theta}\|^{2},\quad\pi_{k}=\frac{N_{k}}{N}.$ (8)
Note that $\mathcal{L}^{0,\mathcal{C}}_{\lambda}$ is equivalent to the standard loss $\operatorname{\mathcal{L}}_{\lambda}$ in the situation of degenerate data, i.e., when the data within any family of $\mathcal{C}$ coincides with its centroid. Moreover, from the previous discussion and Proposition˜3.1, we know that the gradient with respect to any of the last LlL-l layers of the $l$-th centroid-based loss has rank at most $K$. In the following lemma (proved in Section˜A.2), an easy computation allows us to compare the gradients of the original loss and the centroid-based loss:
Lemma 3.3 (Centroid-based gradients approximate full gradients for well-clustered data).
Fix a partition $\mathcal{C}=\{\mathcal{C}_{k}\}_{k=1}^{K}$ and a layer $l$ with weight matrix $W_{l}$. Let the function $\operatorname{\ell}(f_{j}(\mathit{\Theta}_{j},z),y)$ be $C^{1}$ in $\mathit{\Theta}_{j}$ and $\nabla_{W_{l}}\operatorname{\ell}(f_{j}(\mathit{\Theta}_{j},z),y)$ be $C^{2}$ in $(z,y)$ for any $j<l$. Then, $\forall j<l$
$\begin{array}[]{lr}\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}(\mathit{\Theta})=\nabla_{W_{l}}\operatorname{\mathcal{L}}_{0}^{j,\mathcal{C}}(\mathit{\Theta})+R(\mathit{\Theta}),\quad\|R(\mathit{\Theta})\|\leq M_{l,j}(\mathit{\Theta},\mathcal{C})\operatorname{\mathrm{WCSS}}_{j}(\mathcal{C})\end{array}$ (9)
where ${M_{l,j}(\mathit{\Theta},\mathcal{C}):=\sup_{\begin{subarray}{c}k=1,\dots,K\\ (z,y)\in\mathcal{H}^{j}_{k}\end{subarray}}\|\nabla^{2}_{(z,y)}\nabla_{W_{l}}\operatorname{\ell}(f_{j}(\mathit{\Theta}_{j},z),y)\|}$ and $\mathcal{H}^{j}_{k}:=\mathrm{ConvHull}\{(z_{i}^{j},y_{i})\}_{i\in\mathcal{C}_{k}}$.
As a consequence of the above lemma, for any stationary point, we can control the distance of the $l$-th layer from the set of matrices having rank less than or equal to $K$ in terms of the $K$-th total cluster variation of the first l1l-1 intermediate outputs. Before stating our main result, we define the set of $d_{1}\times d_{2}$ matrices having rank equal to $K$, $\mathcal{M}_{K}^{d_{1},d_{2}}:=\{A\in\mathbb{R}^{d_{1}\times d_{2}}|\;\operatorname{rank}(A)=K\}$. We will omit the dimension-related superscripts when clear from the context.
Theorem 3.4 (Small within-class variability yields low-rank bias).
Assume $\mathit{\Theta}^{*}$ to be a stationary point of $\operatorname{\mathcal{L}}_{\lambda}(\mathit{\Theta})$. Then, for any $l=1,\dots,L$ and $K\leq N$,
$\min_{Z\in\cup_{r=1}^{K}\mathcal{M}_{r}}\|W^{*}_{l}-Z\|\leq\min_{j=1,\dots,l-1}\frac{M_{l,j}(\mathit{\Theta}^{*})}{\lambda}\mathrm{TCV}_{K,j},$
where $M_{l,j}(\mathit{\Theta}^{*}):=\!\!\!\!\!\displaystyle{\sup_{(\alpha,\beta)\in\mathcal{H}^{j}}\!\!\|\nabla_{(z_{j},y)}^{2}\nabla_{W_{l}}\\ \operatorname{\ell}(f_{j}(\mathit{\Theta}^{*}_{j},z^{j}),\alpha,\beta)\|}$ and $\mathcal{H}^{j}:=\mathrm{ConvHull}\{(z^{j}_{i},y_{i})\}_{i=1}^{N}$.
Theorem˜3.4 shows that an intermediate neural collapse combined with weight decay regularization leads to approximately low-rank network layers, where the numerical ranks effectively depend on the $\operatorname{\mathrm{TCV}}$ of intermediate layers. Its proof can be found in Section˜A.3. These considerations are directly connected with the notion of deep (or intermediate) neural collapse sukenik2024deep ; papyan2020traces ; rangamani2023feature , whose first characterizing property is recalled below:
Definition 3.5 (Deep Neural Collapse 1).
We say that a neural network $f$ trained on a dataset $\mathcal{X},\mathcal{Y}$ satisfies $\operatorname{\mathrm{DNC1}}$ at layer $l$ if $\operatorname{\mathrm{TCV}}_{K,l}=0$ for $K=c$ and the minimizer in (7) is given by the partition corresponding to the classes.
Notice that the vanishing property of $\operatorname{\mathrm{TCV}}$ in intermediate layers ($\operatorname{\mathrm{DNC1}}$) has already been studied for simplified models in, e.g., Poggio_hidden_nc ; sukenik2024deep ; sukenik2024neural ; jacot2025wide . We therefore notice here the special role of $\operatorname{\mathrm{DNC1}}$ with respect to the other $\mathrm{DNC}$ properties commonly studied, such as the ones presented in sukenik2024deep : the first property alone implies that the stationary points in parameter space must have low-rank, as predicted by Theorem˜3.4.
In the next section, we will prove that, in a fairly general setting, $\operatorname{\mathrm{DNC1}}$ is optimal for all intermediate layers: this means that under a set of natural constraints, a zero total cluster variation for all intermediate layers is globally optimal in the small $\lambda$ regime. This shows, in combination with Theorem˜3.4, that all global minima in this regime are low rank. Moreover, we will show that for almost every initial condition, there is essentially no loss barrier to configurations with $\operatorname{\mathrm{TCV}}=0$ for all $l=1,\dots,L$, shedding new light on the emergence deep neural collapse and, in turn, on the emergence of low-rank bias.
4 Vanishing $\operatorname{\mathrm{TCV}}$: global optimality of deep-neural collapse
We start by defining the representation cost functional for a biasless network with $L^{2}$ regularization in all layers but the first one:
$\mathscr{L}(\mathit{\Theta}):=\begin{cases}+\infty,\quad\text{if}\,\,f(\mathit{\Theta};X)\neq Y,\\ \frac{1}{2}\sum_{l=1}^{L-1}\|W_{l+1}\|^{2},\quad\text{otherwise}.\end{cases}$ (10)
This functional has been studied in a variety of different settings, both linear Ranzato_2021 and nonlinear jacot2023bottleneck ; jacot2023implicit . While in the deep linear case the explicit form of the minima of $\mathscr{L}$ are known to be Shatten quasi-norms Ranzato_2021 ; ongie2022role , in the general case a full characterization is still missing. In (jacot2023implicit, ; jacot2023bottleneck, ), the authors were able to characterize, in the large depth regime, how the minimal value essentially depends on the data $y=f^{*}(x)$ one is trying to interpolate, essentially showing that the first order term scales with $L$ and it is a function bounded between the Jacobian and the bottleneck rank of the function $f^{*}$.
The general intuition is that $\mathscr{L}(\mathit{\Theta})$ represents the loss of interpolating networks in the small $\lambda$ regime, in fact from Scagliotti_2022 , we have that $\lambda^{-1}\mathcal{L}_{\lambda}\underset{\lambda\to 0}{\to}\mathscr{L}(\mathit{\Theta})$ in the $\Gamma$-convergence sense (Definition˜A.1), meaning that minimizers of $\mathcal{L}_{\lambda}$ converge to minimizers of $\mathscr{L}$ as $\lambda$ goes to zero, see Proposition˜A.2. Now, using the observation done in jacot2022featurelearningl2regularizeddnns , we can reparameterize the problem using minimal norm feature matrices in the following way:
$Z^{l+1}=\sigma_{l+1}(W_{l+1}Z^{l})\iff W_{l+1}=\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+},\,\,\text{s.t.}\,\,\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1}),$
where $Z^{l,+}$ denotes the Moore-Penrose pseudoinverse of the feature matrix $Z^{l}$. Notice that the constraint $\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1})$ is effectively a representability condition on the features, equivalent to $W_{l+1}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1})$. With this new representation, we can study the representation cost as a function of the feature matrices $Z^{l}$, namely:
$\mathscr{L}(Z):=\begin{cases}+\infty,\quad\text{if}\,\,Z^{L}\neq Y,\\ \frac{1}{2}\sum_{l=1}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2},\quad\text{otherwise}.\end{cases}$ (11)
As proven in jacot2022featurelearningl2regularizeddnns , it is not restrictive to study this reparametrized representation cost in the sense that the local minimizers of (10), subject to the representability constraints (I), are in bijection with those of (11), therefore making them equivalent. The key observation is that the objective in (11) is local, i.e., it is a sum of local objectives $\mathscr{L}_{l}=\mathscr{L}_{l}(Z^{l+1},Z^{l})$ depending just on neighbour representations. Therefore, since $\|W_{1}\|^{2}$ is not present in the regularization, $\mathscr{L}$ depends on $Z^{1}$ just through $\mathscr{L}_{1}(Z^{2},Z^{1})=\frac{1}{2}\|\sigma_{2}^{-1}(Z^{2})Z^{1,+}\|^{2}$, which can be minimized explicitly with respect to $Z^{1}$, as a function of the next representation $Z^{2}$. For this reason, the lack of the first layer in the regularization term is the key to obtaining the results in the following section, as it makes it possible to study recursively the structure of minima layer by layer. To study it, we consider the additional hypothesis that $Y$ is in one-hot encoding form, i.e. $y_{j}=e_{i_{j}}$ for any $j$, where $\{e_{i}\}$ is the canonical basis of $\mathbb{\mathbb{R}}^{c}$. We also point out that, without considering additional constraints, the minima of $\mathscr{L}$ may not exist. To avoid these situations, we study the minima of $\mathscr{L}$ subject to additional constraints:
Definition 4.1 (Constraints).
Consider the set $\mathcal{S}$ consisting of all networks $Z=(Z^{1},\dots,Z^{L})\in\mathbb{R}^{n_{1}\times N}\times\dots\times\mathbb{R}^{n_{L}\times N}$ that satisfy the following conditions (I)-(II)-(III) for all $l=1,\dots,L-1$:
  1. (I)
    $\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1})$,
  2. (II)
    $s_{i}(Z^{l})\leq Cs_{i}(\sigma_{l+1}^{-1}(Z^{l+1})),\quad\forall i=1,\dots,r_{l}$,
  3. (III)
    $\operatorname{rank}(\sigma_{l}^{-1}(Z^{l}))\geq\operatorname{rank}(Y)$,
where $C>0$ is arbitrary, $r_{l}$ is the rank of $\sigma_{l+1}^{-1}(Z^{l+1})$, $s_{i}(A)$ denotes the $i$-th singular value of a matrix $A$ and $\sigma_{2},\dots,\sigma_{L}$ is a set of entrywise omeomorphisms with $\sigma_{l}(x)\neq 0$ for all $x\neq 0$.
We highlight here that constraint (II) corresponds to a constraint just on the nonzero singular values of the matrices $Z^{l}$ and $\sigma_{l+1}^{-1}(Z^{l+1})$. Therefore, (II) allows the case in which all feature matrices $Z^{l}$ for $l=1,\dots,L-1$ are full-rank. Moreover, $\mathcal{S}$ is bounded but not compact, making it nontrivial to prove the existence of a minimizer. The following main result on the minimizers of a constrained version of (11) holds:
Theorem 4.2 ($\operatorname{\mathrm{DNC1}}$ is optimal for constrained representation cost).
The global minimizers of the optimization problem
$\min_{Z\in\mathcal{S}}\mathscr{L}(Z)$
satisfy $\operatorname{\mathrm{DNC1}}$ for all intermediate layers $l=1,\dots,L$. In particular, any minimum point $\bar{Z}$ satisfies
$\bar{Z}^{l}\propto O_{l}\sigma_{l+1}^{-1}(\bar{Z}^{l+1}),\quad O_{l}\in\mathbb{R}^{n_{l}\times n_{l+1}},\quad O_{l}^{\top}O_{l}=I,\quad\forall\,l=1,\dots,L-1$
where the proportionality constant is $C$ in Definition˜4.1.
A rigorous formulation and proof of Theorem˜4.2 can be found in Appendix˜B.
Theorem˜4.2 can be reinterpreted as follows: if we allow the architecture to collapse already at the first layer, that is, if the kernel of $W_{1}$ can be large enough (i.e., $n_{1}\geq N$), then, under (I)(III), it is optimal to collapse all intermediate layers. Notice that the first set of constraints (I) has a crucial role in terms of the representability of the features, and it effectively couples the representations of different layers, as discussed above. The second constraint (II) can be directly related to the boundedness of the nonlinear activations, and it represents exactly the situation for diagonal feature matrices, for which the nonlinearity acts directly on the singular values. The necessity of constraint (II) relies in making it possible to study the alignment of the optimal basis, avoiding any possible singular value escaping to infinity. It is interesting to notice that while in (II) $C$ is an arbitrary positive constant, we observe numerically in Section˜5 that a perfect fit happens for $C=1$, reminiscent of a norm balancing condition on all the layers. The third set (III) is introduced here as it arises naturally in the proof of the theorem as a consequence of (II), and it allows us to rule out the presence of global minima of lower rank. This is because, as we will show in the proof in Appendix˜B, all global minima satisfy the equality constraint. We notice that, as a byproduct of Theorem˜4.2, we know that if we get rid of constraint (III), then all global minimizers either satisfy $\operatorname{\mathrm{DNC1}}$ or some layer embedding has rank strictly smaller than $K$. We expect these minimizers to be exceptional in some sense, as they were proven to be optimal in a very particular setting in sukenik2024neural . Morever, we point out that also the singular values structure is predicted recursively by Theorem˜4.2, which we will numerically show also in Section˜5.
We emphasize that Theorem˜4.2 can be extended to residual networks under a weaker set of assumptions. Indeed, in that case, we can get rid of all constraints, except the natural one in (I). Due to space limitations, we present this result along with its proof in Appendix˜C.
In the case of feedforward networks, with a stronger request on the intermediate widths, we have the following additional main result, proving that for almost every interpolating set of parameters, there exists a continuous path in parameter space along which the loss decreases, finally converging to globally optimal points, which satisfy $\operatorname{\mathrm{DNC1}}$ for Theorem˜4.2.
Theorem 4.3 ($\operatorname{\mathrm{DNC1}}$ is reachable for almost every initial condition).
Assume the intermediate widths in $\mathcal{S}$ satisfy the condition $K\leq n_{L}=n_{L-1}=\dots=n_{1}=N$ and that the $\sigma_{l}$’s are analytic. Let $Z(0)\in\mathcal{S}$ be a random initial interpolating condition, i.e., such that $Z^{L}(0)=Y$. Then, almost surely there exists a continuous path $\gamma(t)$ with $\gamma(0)=Z(0)$, $\gamma(1)$ global minimum of $\mathscr{L}$, and such that $\mathscr{L}\circ\gamma$ is continuous Lebesgue-almost everywhere (except for a finite set of times) and for all continuity times $s\geq t$ we have
$\mathscr{L}(\gamma(t))\geq\mathscr{L}(\gamma(s)).$
Moreover, all discontinuities $t_{1},\dots,t_{m}$ are removable and satisfy
$\underset{t\to t_{i}^{-}}{\lim}\mathscr{L}(\gamma(t))=\underset{t\to t_{i}^{+}}{\lim}\mathscr{L}(\gamma(t))>\mathscr{L}(\gamma(t_{i})),\quad\forall i=1,\dots,m.$
The proof of Theorem˜4.3 can be found in Section˜B.1. The main takeaway is that, from almost every interpolating initial condition (aside from a set of measure zero), one can construct a continuous path in parameter space along which the loss decreases (up to a finite set of downward jumps) and the network converges to a global minimum characterized by deep neural collapse. We highlight that in any numerical method, one looks at a discretized version of the path $\gamma$, so the “probability" of seeing discontinuity points is zero, no matter how small the learning rate. In particular, no matter how small the stepsize, there always exists a sequence of points that decreases the loss and reaches $\operatorname{\mathrm{DNC1}}$ configurations. This result is exactly formalized in Corollary˜B.4 and Remark˜B.5. In particular, a consequence of Theorem˜4.3 is that $\inf_{\gamma\in\Gamma}\sup_{t\in[0,1]}\mathscr{L}(\gamma(t))\leq\mathscr{L}(Z(0))$ where $\Gamma:=\{\gamma\in\mathcal{C}^{0}([0,1];\mathcal{S})\,|\,:\gamma(0)=Z(0),\gamma(1)\,\text{satisfies}\,\,\operatorname{\mathrm{DNC1}}\,\,\forall l=1,\dots,L\}$.
A second way to interpret Theorem˜4.3 and Corollary˜B.4 is that the loss landscape of $\mathscr{L}$, despite being highly nonconvex, is well behaved in the sense that there are essentially no loss barriers from almost all interpolating points to global minima. This result, being independent of the specific temporal dynamics imposed during training, provides theoretical support for the empirical “universality” of neural collapse observed in classification tasks when using sufficiently wide architectures Poggio_hidden_nc and possibly gives an hint on why it is observed towards the end of training papyan2020prevalence ; Poggio_hidden_nc (i.e., in the interpolating phase, where the landscape is well behaved).
Moreover, the phenomenon is “unidirectional”, progressing from high to low rank as demonstrated in the proof, which is reminiscent of the behaviour observed in the deep linear case in wang2024implicit .
Finally, we point out that these results are recovered as the singular limit in the regime where an arbitrarily small penalization term for the first layer $\|W_{1}\|^{2}=\|\sigma_{1}^{-1}(Z^{1})X^{+}\|^{2}$ is included. A natural question arises: are the highly structured solutions given by Theorem˜4.2 preserved when $\|W_{1}\|^{2}$ is included? To answer this question about stability, we consider the perturbed loss
$\mathscr{L}_{\lambda}(Z)=\begin{cases}+\infty,\quad\text{if}\,\,\ Z^{L}\neq Y,\\ \frac{1}{2}\sum_{l=1}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{1}{\lambda}\|\sigma_{1}^{-1}(Z^{1})X^{+}\|^{2},\quad\text{otherwise},\end{cases}$ (12)
defined on the set $\mathcal{S}$, see (4.1), and show that, as $\lambda\to\infty$, the global minima of $\mathscr{L}_{\lambda}$ converge to global minima of $\mathscr{L}$. Note that varying the value of $\lambda$ is equivalent to rescaling the dataset $X$ by a factor $\lambda$.
Theorem 4.4 (Stability of the $\operatorname{\mathrm{DNC1}}$ global minima).
The functionals $\mathscr{L}_{\lambda}:\mathcal{S}\to\mathbb{R}$ converge to $\mathscr{L}:\mathcal{S}\to\mathbb{R}$ as $\lambda\to+\infty$, in the $\Gamma$-convergence sense.
Moreover, as a classical consequence of $\Gamma$-convergence, if $Z^{*}\in\mathcal{S}$ is an accumulation point for a sequence of minimizers $Z_{n}\in\arg\min_{Z\in\mathcal{S}}\mathscr{L}_{\lambda_{n}}(Z)$ with $\lambda_{n}\to+\infty$, then $\mathscr{L}(Z^{*})=\min_{Z\in\mathcal{S}}\mathscr{L}(Z)$. The proof is included in Section˜B.1. We also highlight here that the existence of an accumulation point $Z^{*}$ is not immediate, since $\mathcal{S}$ is not compact.
The interpretation of Theorem˜4.4 is twofold. First, the optimality of $\operatorname{\mathrm{DNC1}}$ configurations in Theorems˜4.2 and C.2 are stable with respect to the presence of regularization on $\|W_{1}\|^{2}$, therefore, we expect configurations close to $\operatorname{\mathrm{DNC1}}$ for all layers even when the first weight decay is present. Moreover, the regime in which one may expect configurations to be close to $\operatorname{\mathrm{DNC1}}$ ones with the first weight decay included is when $\|X\|\gg\|Y\|.$ Alternatively, another possible interpretation of $\lambda$ in Theorem˜4.4 is as the depth of the network. In fact, the minimizers of Equation˜12 and those of Equation˜12 divided by $L$ are the same. The first term, being bounded on the minimizers, is vanishing for $L\to+\infty$. Therefore, Theorem˜4.4 is essentially telling that one should expect the minimizer structure as the one predicted in Theorem˜4.2 for large depth networks. Furthermore, we remark that the regularization term due to any finite number of initial layers vanishes in the large depth limit, making the results of Theorem˜4.2 hold also for less restrictive conditions on the widths, but from the first layer $l^{*}$ which can represent a large enough region of the full space $\mathbb{R}^{n_{l^{*}}\times N}$ (the one containing the minimizers predicted in Theorem˜4.2). A more precise discussion of this can be found in Remark˜B.6. An equivalent stability result holds even for residual networks, as discussed in Remark˜C.3.
5 Experimental evaluation
5.1 Qualitative behaviour of Theorem˜3.4
In this section, we show in a simple setting the qualitative dependency of $\min_{Z\in\cup_{r=1}^{K}\mathcal{M}_{r}}\|W^{*}_{l}-Z\|$ on the regularization parameter $\lambda$ and on the TCV of the dataset. In particular, fixed a set of vectors $\mu_{1},\dots,\mu_{K}$, we consider a random Gaussian mixture dataset $X\in\mathbb{R}^{n\times N}$ with $x_{i}\sim N(\mu_{y_{i}},\sigma^{2})$, and a $Y\in\mathbb{R}^{K\times N}$ one-hot encoding matrix with $K=10$ classes. In this way, we know that $\operatorname{\mathrm{TCV}}_{K,0}\sim O(\sigma^{2})$, and we expect, from Theorem˜3.4, the ranks to approach $K=10$ as $\sigma^{2}\to 0$. Then, for each $(\lambda,\sigma)$ in a grid of values, we sampled $X,Y$ according to the distribution described above, and we trained a four-layer fully connected neural network with the loss defined in Equation˜3. In Figure˜1 we report the results of this experiment. We can observe that consistently, for each $\lambda$ in our grid, we have that all matrices are essentially of rank $K$ for $\sigma$ small enough. Moreover, we also observe the decreasing behavior with respect to $\lambda$ consistently with Theorem˜3.4.
Figure 1: For each couple $(\lambda,\sigma)$ we report the relative distance $\sum_{j>K}s_{j}^{2}/\sum_{j}s_{j}^{2}$ of trained weight matrices from the closest rank-$K=10$. Here, $s_{j}$ are the singular values of each weight matrix.
5.2 Convergence to the optimal solution of Theorem˜4.3
We start by illustrating and numerically validating what we claimed in Theorems˜4.3 and 4.4. In particular, by setting $\delta=1/\lambda$ in (12) we can show that for $\delta\to 0$, the singular values of all the intermediate post-activation features are the ones described in Theorem˜4.2. To set up the experiment, we subsampled the MNIST dataset deng2012mnist with $N=200$ examples, and we considered a $5$ layer feedforward network with all $\sigma_{l}=\tanh$ activations and widths equal to $n_{1}=\dots=n_{4}=N=200,n_{5}=K=10$. With $\delta$, the loss becomes $\mathcal{L}_{\lambda,\delta}=\frac{1}{2}\|f_{\mathit{\Theta}}(X)-Y\|^{2}+\frac{\lambda}{2}\sum_{l=2}^{L}\|W_{l}\|^{2}+\frac{\lambda\delta}{2}\|W_{1}\|^{2}+\frac{\lambda^{\prime}}{2}\sum_{l=1}^{L}\|Z^{l}\|^{2}$, where the last regularization term can be thought of as a weaker way of imposing the second set of constraints in Theorem˜4.3. In particular, for this experiment we considered $\lambda=\lambda^{\prime}=10^{-5}$.
We trained the network until convergence (way past interpolation, with a final data loss lower than $10^{-6}$ in all examples) for different values of $\delta$, and we compared the actual singular values of the feature matrices after training with the ones predicted by Theorem˜4.2. The results are shown in Figure˜2. As we can notice, the singular values are converging to the ones predicted by Theorem˜4.2, and the convergence with respect to $\delta$ seems to be fast. In fact, already for moderate sizes of $\delta$ (e.g., $\delta=2$), the singular values are close to the predicted ones. Additional results are included in Figure˜4.
$\delta=10$ $\delta=5$ $\delta=2$ $\delta=0.01$
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Singular values of trained networks (full line) against those predicted by Theorem˜4.2 (dots) for different values of $\delta=\frac{1}{\lambda}\in\{10,5,2,0.1\}$. The singular values are converging to the predicted ones as $\delta\to 0$, as shown by Theorem˜4.4 in combination with Theorem˜4.2.
5.3 Convergence to $\operatorname{\mathrm{DNC1}}$ configurations
In this experiment, we will showcase the convergence of $\operatorname{\mathrm{TCV}}$ to zero of all intermediate layers in the setting of Theorem˜4.2. The experimental setting is the same as in Section˜5.2, but with $L=10$ layers and $\delta=0$ fixed. In Figure˜3 (rightmost figure), we plot for a single run the $\operatorname{\mathrm{TCV}}$ of all intermediate post-activation embeddings. As we can observe from this first plot, the $\operatorname{\mathrm{TCV}}$ of all intermediate embeddings goes essentially to machine precision as expected from Theorems˜4.2 and 4.3. In Figure˜3 (first three figures from the left), we instead plot the average $\operatorname{\mathrm{TCV}}$ across all intermediate layers, for $100$ different random unitary initializations, to avoid slow convergence caused by the depth. As we can see, the $\operatorname{\mathrm{TCV}}$ consistently goes to zero in all runs. Regarding the numerical ranks (up to $10^{-5}$ relative error), we observe that they are all close to the optimal ones ($K=10$).
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: First three plots from the left: average of different quantities on all intermediate layers and for $100$ different random initializations. Shadows represent the area included between the minimal and maximal values observed. Each optimization step represents 100 epochs. Right figure: TCV of each single layer during training.
6 Conclusions
We proposed a theoretical investigation on the emergence of low-rank weight matrices and deep-neural collapse for a general class of neural networks. In particular, we established a quantitative relationship between deep neural collapse and the rank collapse of the weight matrices. In addition, we proved that for a general class of models, by not regularizing the first layer, $\operatorname{\mathrm{DNC1}}$ is optimal for all intermediate representations under a set of natural constraints, and that essentially there is no loss barrier between the interpolating global minima and $\operatorname{\mathrm{DNC1}}$ configurations. Lastly, we showed that the $\operatorname{\mathrm{DNC1}}$ configurations are stable and are preserved even when the first weight decay is included. Future investigations will go in the direction of relaxing some constraints and trying to adapt this approach to attention-based networks and GNNs. Moreover, we expect the same tools to be of interest when studying the representation learning of DNNs in general, which is also planned for future investigation.
Appendix A Useful definitions and results
In this section, we will introduce some useful definitions and fact which we will need to prove the results in Appendix˜B. We start recalling the definition and some basic facts about $\Gamma$-convergence of functionals (see [6]).
Definition A.1.
($\Gamma$-convergence) Let $S$ be a locally numerable topological space and let $F_{n}:S\to\mathbb{R}$ be a sequence of functionals and $F:S\to\mathbb{R}$ be another functional. $F_{n}$ is said to $\Gamma$-converge to $F$ if the following hold:
  • (i)
    For any sequence $(x_{n})\subset S$ with $x_{n}\to x$, we have
    $F(x)\leq\liminf_{n\to\infty}F_{n}(x_{n}).$
  • (ii)
    (Recovery sequence condition). For any $x\in S$, there exists a sequence $x_{n}\to x$ such that
    $F(x)\geq\limsup_{n\to\infty}F_{n}(x_{n}).$
To build some intuition, the first condition means that asymptotically, $F$ is a common lower bound for the sequence of $F_{n}$, at least pointwise. The second condition is a condition of recovery of the lower bound, implying, in a certain sense, optimality of the lower bound. This definition can be useful when it is simpler to characterize the minima of $F_{n}$ than those of $F$, or viceversa. In our case, we can characterize the minima of $F$, but not those of $F_{n}$. However, thanks to $\Gamma$-convergence, we can establish that the minima of $F_{n}$ are going be close to those of $F$ in suitable scenarios. In particular, the following well known result holds:
Proposition A.2.
[6] Let $F_{n}\to F$ in the sense of $\Gamma$-convergence and let $S$ be sequentially compact. Then, for every sequence $x_{n}\in\arg\min_{y\in S}F_{n}(y)$ and every limit point $x$ of $(x_{n})$, we have
$x\in\arg\min_{y}F(y).$
The next lemma discusses the zeros of single-variable analytic functions, showing that their cardinality is always finite on compact sets, which we will need later.
Lemma A.3 (Finite number of zero of real analytic functions on compact intervals, [19]).
Consider a nonzero real analytic function $f:[0,1]\to\mathbb{R}$. Then $f$ has a finite number of zeros in $[0,1]$.
Proof.
First, we show that zeros are isolated. Let $t_{0}\in[0,1]$ be a zero of $f$. Since $f$ is real analytic, then we can write
$f(t)=\sum_{k=0}^{+\infty}a_{k}(t-t_{0})^{k},\quad|t-t_{0}|\leq\rho$
since $f\neq 0$, there exists $m$ such that $a_{m}\neq 0$. Therefore,
$f(t)=(t-t_{0})^{m}h(t),\quad h(t):=\sum_{k=0}^{+\infty}a_{m+k}(t-t_{0})^{k},\quad h(t_{0})=a_{m}\neq 0$
Therefore, by continuity of $h$ there exists a neighborhood of $t_{0}$ of radius $\rho^{\prime}$ such that $h(t)\neq 0$ for all $|t-t_{0}|<\rho^{\prime}$. Then in this interval we have $f(t)=0\iff t=t_{0}$, proving that $t_{0}$ is isolated. Now, assume $f$ has infinitely many zeros $(t_{k})$ in $[0,1]$, then by compactness there exists a convergent subsequence $t_{k_{j}}\to\bar{t}\in[0,1]$. But then, $\bar{t}$ would be an accumulation point, and this is only possible if $f$ is zero in an interval around $\bar{t}$. Since $f$ is analytic, this would force $f\equiv 0$, which is a contradiction. ∎
Proposition A.4 (Lower semicontinuity of functions defined through infimum).
Let $X,Y$ be two finite-dimensional normed spaces, $f:X\times Y\to\mathbb{R}$ and $g:X\times Y\to\mathbb{R}^{m}$ be continuous functions, and $\Omega(x)=\{y\in Y\,|g(x,y)=0\,\}$. Moreover, assume that $f(x,y)$ is coercive with respect to $y$, uniformly in $x$. Then, the function
$F(x)=\inf_{y\in\Omega(x)}f(x,y)$
is lower semicontinuous.
Proof.
To prove the lower semicontinuity of $F(x)$ we need to show that, for every sequence $x_{k}\to\bar{x}$, we have
$\liminf_{k}F(x_{k})\geq F(\bar{x}).$
Let $(x_{k_{j}})$ be a subsequence of $(x_{k})$ such that $\exists\lim_{j}F(x_{k_{j}})=\liminf_{k}F(x_{k})$ Then by definition of infimum, for every $x_{k}$ we can find $y_{k}\in\Omega(x_{k})$ such that $F(x_{k})\leq f(x_{k},y_{k})\leq F(x_{k})+\frac{1}{k}$. Then we have two cases:
  • (i)
    If the sequence $(y_{k_{j}})$ is unbounded, up to considering a subsequence, we can assume that $\|y_{k_{j}}\|\to\infty$. Then, for the uniform coercivity of $f$ we have
    $f(x_{k_{j}},y_{k_{j}})\to\infty$
    and, therefore,
    $\liminf_{k}F(x_{k})=\lim_{j}F(x_{k_{j}})\geq\lim_{j}\left(f(x_{k_{j}},y_{k_{j}})-\frac{1}{{k_{j}}}\right)=+\infty\geq F(\bar{x}).$
  • (ii)
    If the sequence $(y_{k_{j}})$ is bounded, up to considering a subsequence, we can assume $y_{k_{j}}\to\bar{y}$. Since $y_{k_{j}}\in\Omega(x_{k_{j}})$ for all $j$, by the continuity of $g$, $0\equiv g(x_{k_{j}},y_{k_{j}})\to g(\bar{x},\bar{y})$. Therefore, $\bar{y}\in\Omega(\bar{x})$. Similarly, the continuity of $f$ gives
    $f(x_{k_{j}},y_{k_{j}})\to f(\bar{x},\bar{y}).$
    Therefore, we see that
    $\liminf_{k}F(x_{k})=\lim_{j}F(x_{k_{j}})\geq\lim_{j}\left(f(x_{k_{j}},y_{k_{j}})-\frac{1}{k_{j}}\right)=f(\bar{x},\bar{y})\geq\inf_{y\in\Omega(\bar{x})}f(\bar{x},y)=F(\bar{x}),$
    which is the required inequality.
This concludes the proof. ∎
Remark A.5.
We will use this last result in Section˜B.1 to prove the lower semicontinuity of $F(A,B)=\|AB^{+}\|^{2}$. In fact, by defining
$x=(A,B)\in X=\{(A,B)\in\mathbb{R}^{n\times N}\times\mathbb{R}^{m\times N}\,|\,AB^{+}B=A\}\quad\text{ and }\quad y=W\in\mathbb{R}^{n\times m}$
we can define g(x,y)=WBAg(x,y)=WB-A, $f(x,y)=\|W\|^{2}$ and
$F(x)=\inf_{y\in\Omega(x)}f(x,y),$
recalling that $\Omega(x)=\{y=W\;|\;g(x,y)=WB-A=0\}$ and using a matrix version of [10, Section 5.5.1]. In particular we are exactly in the setting of Proposition˜A.4, proving lower semicontinuity of $F(A,B)=\|AB^{+}\|^{2}$.
Lemma A.6 (Blowup of $\mathscr{L}$ outside the natural constraint).
Consider a fixed matrix $A\in\mathbb{R}^{n\times N}$ with $N\geq n\geq\operatorname{rank}(A)=r$ and the matrix functions $f,g:\mathbb{R}^{m\times N}\to\mathbb{R}$ defined as $g(B)=\|AB^{+}B-A\|^{2}$ and $f(B)=\|AB^{+}\|$ with $m\geq n$. Let $B_{k}\rightarrow\bar{B}$ with $g(B_{k})=0$ for every $k$ and $g(\bar{B})\neq 0$. Then, $\lim_{k}f(B_{k})=+\infty$.
Proof.
To prove it, let $U_{k}\Sigma_{k}V_{k}^{T}$ be the reduced SVD decomposition of $B_{k}$ and $\bar{U}\bar{\Sigma}\bar{V}^{T}$ the one of $\bar{B}$. Then by assumption we have $A=A\bar{V}\bar{V}^{T}+A(I-\bar{V}\bar{V}^{T})$ with the second term on the right hand side being different from zero because $g(\bar{B})\neq 0$. Additionally, w.l.o.g., for any $k$ we can write $V_{k}=[V_{k,1},V_{k,2}]$, $\Sigma_{k}=\mathrm{diag}[\Sigma_{k,1},\Sigma_{k,2}]$ and $U_{k}=[U_{k,1},U_{k,2}]$ where $U_{k,1},\Sigma_{k,1},V_{k,1}]\rightarrow[\bar{U},\bar{\Sigma},\bar{V}]$, $I-V_{k,1}{V_{k,1}}^{T}\rightarrow(I-\bar{V}\bar{V}^{T})$ and $\Sigma_{k,2}\rightarrow 0$. Then
$f(B_{k})$ $=\|AV_{k}^{1}\Sigma_{k,1}^{-1}U_{k,1}^{T}+AV_{k,2}\Sigma_{k,2}^{-1}U_{k,2}^{T}\|\geq\|AV_{k,2}\Sigma_{k,2}^{-1}\|-\|AV_{k,1}\Sigma_{k,1}^{-1}\|$ (13)
$\geq\min_{i}{(\Sigma_{k,2})_{ii}^{-1}}\|AV_{k,2}\|-\|AV_{k,1}\Sigma_{k,1}^{-1}\|\rightarrow+\infty-\|A\bar{V}\bar{\Sigma}^{-1}\|=+\infty,$
where we have used that the singular values are all nonnegative, i.e., $\Sigma_{k_{ii}}\geq 0$ for all $i$ and $k$ with ${\Sigma_{k,2}}_{ii}\rightarrow 0$ for any i, that $\lim_{k}AV_{k,2}=A(I-\bar{V})\neq 0$ and the equalities $\|AV_{k,1}\Sigma_{k,1}^{-1}U_{k,1}^{T}\|^{2}=Tr(AV_{k,1}\Sigma_{k,1}^{-1}U_{k,1}^{T}U_{k,1}\Sigma_{k,1}^{-1}V_{k,1}^{T}A)=Tr(AV_{k,1}\Sigma_{k,1}^{-1}\Sigma_{k,1}^{-1}V_{k,1}^{T}A)=\|AV_{k,1}\Sigma_{k,1}^{-1}\|^{2}$ and analogously $\|AV_{k,2}\Sigma_{k,2}^{-1}U_{k,2}^{T}\|=\|AV_{k,2}\Sigma_{k,2}^{-1}\|$. ∎
A.1 Proof of Proposition 3.1
The main step of the proof is to show that the gradient with respect to a single data point produces matrices of rank at most 1. One then concludes by observing that the sum of $K$ terms of rank at most $1$ has rank at most $K$. Given the structure of the loss function that depends only on the output of the model $f(\mathit{\Theta},x)$, if we use the chain rule and the recursive definition of $f(\mathit{\Theta},x)$ in (1), we obtain
$\nabla_{W_{l}}\operatorname{\ell}\big{(}f(\mathit{\Theta},x),y\big{)}$ $=\nabla_{W_{l}}\operatorname{\ell}\big{(}f_{l}(\mathit{\Theta}_{l},z^{l}(x)),y\big{)}=\frac{\partial f(\mathit{\Theta},x)}{\partial W_{l}}^{\top}\nabla_{f}\operatorname{\ell}\big{(}f(\mathit{\Theta},x),y\big{)}$
$=\frac{\partial z^{l}(x)}{\partial W_{l}}^{\top}\frac{\partial f_{l}(\mathit{\Theta}_{l},z^{l}(x))}{\partial z^{l}}^{\top}\nabla_{f}\operatorname{\ell}\big{(}f(\mathit{\Theta},x),y\big{)}.$
Now we use the assumption that $z^{l}=\sigma(W_{l}z^{l-1}+b_{l})=\sigma(a^{l})$, where we have introduced the auxiliary variable $a^{l}:=W_{l}z^{l-1}+b_{l}$. This finally leads to
$\nabla_{W_{l}}\operatorname{\ell}\big{(}f(\mathit{\Theta},x),y\big{)}$ $=\big{(}z^{l-1}(x)\otimes I\big{)}\frac{\partial z^{L}(x)}{\partial a^{l}}^{\top}\nabla_{z^{L}}\operatorname{\ell}\big{(}z^{L}(x),y\big{)}$
$=\frac{\partial f_{l}(\mathit{\Theta}_{l},\sigma(a^{l}(x)))}{\partial a^{l}}^{\top}\Bigl{(}\nabla_{f}\operatorname{\ell}\big{(}f(\mathit{\Theta},x),y\big{)}\cdot z^{l-1}(x)^{\top}\Bigr{)}.$
By noticing that the last term $\nabla_{f}\operatorname{\ell}\big{(}f(\mathit{\Theta},x),y\big{)}z^{l}(x)^{\top}$ is a matrix of rank at most 1, we can conclude by submultiplicativity of the rank. $\square$
A.2 Proof of Lemma 3.3
For simplicity, we will omit the superscript $j$ from $z^{j}_{i}$. We start by considering the Taylor expansion of the gradient $\nabla_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},z),y\big{)}$ centered in $\big{(}\bar{z}_{k},\bar{y}_{k}\big{)}$ for any point $\big{(}z_{i},y_{i}\big{)}$ with $i\in\mathcal{C}_{k}$ and any $k=1,\dots,K$
$\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},z_{i}),y_{i}\big{)}=$ $\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\bar{z}_{k}),\bar{y}_{k}\big{)}$ (14)
$+\operatorname{\nabla}_{(z,y)}\nabla_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\bar{z}_{k}),\bar{y}_{k}\big{)}\big{[}(z_{i},y_{i})-(\bar{z}_{k},\bar{y}_{k})\big{]}+R_{i}^{k}(\mathit{\Theta}),$
where the reminder is given by
$R_{i}^{k}(\mathit{\Theta}):=\frac{1}{2}\operatorname{\nabla}_{(z,y)}^{2}\nabla_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\alpha_{i}^{k})\beta_{i}^{k}\big{)}[(z_{i},y_{i})-(\bar{z}_{k},\bar{y}_{k}),(z_{i},y_{i})-(\bar{z}_{k},\bar{y}_{k})],$ (15)
for some $(\alpha_{i}^{k},\beta_{i}^{k})$ between $(z_{i},y_{i})$ and $(\bar{z}_{k},\bar{y}_{k})$. Therefore, using the linearity of $\operatorname{\nabla}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\bar{z}_{k})\bar{y}_{k}\big{)}[\cdot]$, we have
$\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}^{j}(\mathit{\Theta})=$ $\frac{1}{N}\sum_{j=1}^{N}\nabla_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},z_{j}),y_{j}\big{)}=$
$=$ $\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in\mathcal{C}_{k}}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},z_{i}),y_{i}\big{)}=$
$=$ $\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in\mathcal{C}_{k}}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\bar{z}_{k}),\bar{y}_{k}\big{)}+$
$+\operatorname{\nabla}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\bar{z}_{k}),\bar{y}_{k}\big{)}[(z_{i},y_{i})-(\bar{z}_{k},\bar{y}_{k})]+R_{i}^{k}(\mathit{\Theta})=$
$=$ $\sum_{k=1}^{K}\pi_{k}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\bar{z}_{k}),\bar{y}_{k}\big{)}+$
$+\frac{1}{N}\sum_{k=1}^{K}\operatorname{\nabla}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\bar{z}_{k}),\bar{y}_{k}\big{)}\bigg{[}\underbrace{\sum_{i\in\mathcal{C}_{k}}((z_{i},y_{i})-(\bar{z}_{k},\bar{y}_{k}))}_{=0}\bigg{]}+$
$+\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in\mathcal{C}_{k}}R_{i}^{k}(\mathit{\Theta})=:\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}^{j,\mathcal{C}}(\mathit{\Theta})+R(\mathit{\Theta}),$
where $\pi_{k}=N_{k}/N$. Hence, we obtain
$\|R(\mathit{\Theta})\|\leq$ $\frac{1}{N}\sum_{k=1}^{K}\sum_{i\in\mathcal{C}_{k}}\|R_{i}^{k}(\mathit{\Theta})\|\leq\frac{M_{\operatorname{\ell},j}(\mathit{\Theta})}{2N}\sum_{k=1}^{K}\sum_{i\in\mathcal{C}_{k}}\|(z_{i}^{k},y_{i}^{k})-(\bar{z}_{k},\bar{y}_{k})\|^{2}=\frac{M_{\operatorname{\ell},j}(\mathit{\Theta})}{2}\operatorname{\mathrm{WCSS}}_{j},$
where
$M_{\operatorname{\ell},j}(\mathit{\Theta})=\sup_{k=1,\dots,K}\sup_{i\in\mathcal{C}_{k}}\|\operatorname{\nabla}^{2}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\alpha_{i}^{k}),\beta_{i}^{k}\big{)}\|\leq\sup_{k=1,\dots,K}\sup_{(\alpha,\beta)\in\mathcal{H}_{k}^{j}}\|\operatorname{\nabla}^{2}_{(z,y)}\operatorname{\nabla}_{W_{l}}\operatorname{\ell}\big{(}f_{j}(\mathit{\Theta}_{j},\alpha),\beta\big{)}\|,$ (16)
as desired. $\square$
A.3 Proof of Theorem˜3.4
Observe that from (4), at the stationary point we have
$W^{*}_{l}=-\frac{1}{\lambda}\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}(\mathit{\Theta}^{*}).$ (17)
Now, Lemma 3.3 allows us to control the distance of the gradient in the right-hand side of (17) from $\mathcal{M}_{K}^{d_{l-1},d_{l}}$. In particular, given $\mathcal{C}\in\cup_{r=1}^{K}\mathcal{P}_{r}$, for any $j=1,\dots,l-1$ we have
$\min_{Z\in\cup_{r=1}^{K}\mathcal{M}_{r}}\|W^{*}_{l}-Z\|$ $\leq\frac{\|\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}(\mathit{\Theta}^{*})-\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}^{j,\mathcal{C}}(\mathit{\Theta}^{*})\|}{\lambda}$
$\leq\frac{M_{l,j}(\mathit{\Theta}^{*},\mathcal{C})\operatorname{\mathrm{WCSS}}_{j}(\mathcal{C})}{\lambda},$
where we have used that $\operatorname{\nabla}_{W_{l}}\operatorname{\mathcal{L}}_{0}^{j,\mathcal{C}}(\mathit{\Theta}^{*})/\lambda$ has rank smaller than or equal to the number of families of $\mathcal{C}$. Then, first minimizing over $j=1,\dots,l-1$, second bounding from above $M_{l,j}(\mathit{\Theta}^{*},\mathcal{C})\leq M_{l,j}(\mathit{\Theta}^{*})$ for any $j$, and third minimizing over all the possible partitions $\mathcal{C}\in\cup_{r=1}^{K}\mathcal{P}_{r}$ yield the thesis.
Appendix B Global minima of representation cost for feedforward networks
In this section, we will prove the results presented in Section˜4 together with the necessary lemmas. As we already discussed in Section˜4, the main idea is to study the minima of $\mathscr{L}$ recursively layer by layer, i.e., to study the minima of the local terms $\mathscr{L}_{loc}(Z^{l+1},Z^{l})=\frac{1}{2}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}$ in $Z^{l}$ for $Z^{l+1}$ fixed. We characterize such minimum, under the constraints considered in Theorem˜4.2, in the next technical lemma. This will serve as a basis for recursively proving the main Theorem˜4.2. We start formally introducing the constraints. Consider $A\in\mathbb{R}^{n\times N}$ a fixed matrix playing the role of $\sigma_{l+1}^{-1}(Z^{l+1})$, with $\operatorname{rank}(A)=r$, and $\{s_{i}(A)\}$ its singular values. Then, for some fixed $m\geq n$ and $C>0$ we write
$\mathcal{S}_{loc}(A):=\Bigl{\{}B\in\mathbb{R}^{m\times N}:s_{i}(B)^{2}\leq Cs_{i}(A)^{2}\,\,\forall i=1,\dots,r,\,\,AB^{+}B=A\Bigr{\}}.$ (18)
Lemma B.1 (Optimal single layer for feedforward representation cost with singular value constraint).
Let $A\in\mathbb{R}^{n\times N}$ be a fixed matrix with $\operatorname{rank}(A)=r$, and $\{s_{i}(A)\}$ its singular values and let also $\mathscr{L}_{loc}(B)=\frac{1}{2}\|AB^{+}\|^{2}$. Then, for $n\leq m$ and any fixed constant $C>0$ in the definition of $\mathcal{S}_{loc}(A)$, we have that the solution to the constrained optimization problem $\min_{B\in\mathcal{S}_{loc}(A)}\mathscr{L}_{loc}(B)$ is attained at all points
$B^{*}=OA,\quad O\in\mathbb{R}^{m\times n},\quad O^{\top}O=I.$
In particular, for any optimal $B^{*}$, we have $\mathscr{L}(B^{*})=\operatorname{rank}(A)/2C$. Moreover, for $n=m=N$, and for Lebesgue-almost every $B_{0}\in\mathcal{S}$, there exists a continuous path $\gamma:[0,1]\to\mathbb{R}^{N\times N}$ with $\gamma(0)=B_{0}$, $\gamma(1)\in\underset{B\in\mathcal{S}_{loc}(A)}{\arg\min}\,\mathscr{L}_{loc}(B)$ and $\frac{d}{dt}\mathscr{L}_{loc}(\gamma(t))\leq 0$.
Proof.
First of all, we prove the existence of a minimizer. Observe that whenever a sequence $B_{k}\to\bar{B}$ with $A\bar{B}^{+}\bar{B}\neq A$, then $\mathscr{L}_{loc}(B_{k})\to+\infty$ thanks to Lemma˜A.6. Therefore, we can restrict our attention to sequences $B_{k}\to\bar{B}$ with $A\bar{B}^{+}\bar{B}=A$ (and therefore in particular with $\operatorname{rank}(\bar{B})\geq\operatorname{rank}(A)$).
Since the constraint set $\mathcal{S}_{loc}(A)$ is bounded, any sequence $B_{k}$ for which $\mathscr{L}_{loc}(B_{k})\to\inf_{B\in\mathcal{S}}\mathscr{L}_{loc}(B)$ is also bounded. Therefore, there exists a convergent subsequence $B_{k_{j}}\to\bar{B}$ with $\operatorname{rank}(\bar{B})\geq\operatorname{rank}(A)$. By lower semicontinuity of $\mathscr{L}_{loc}$ (shown in Remark˜A.5) we obtain
$\lim_{j}\mathscr{L}_{loc}(B_{k_{j}})=\liminf_{j}\mathscr{L}_{loc}(B_{k_{j}})=\inf_{B\in\mathcal{S}}\mathscr{L}_{loc}(B)\geq\mathscr{L}_{loc}(\bar{B}).$
Since the singular values $s_{i}(B)$ are continuous functions of $B$, the inequality $s_{i}(B)^{2}\leq Cs_{i}(A)^{2}$ gives a closed set, and therefore $s_{i}(\bar{B})^{2}\leq Cs_{i}(A)^{2}$ . Moreover, we restricted without loss of generality to sequences $(B_{k})$ for which $A\bar{B}^{+}\bar{B}=A$ and therefore we have $\bar{B}\in S_{loc}$. Combining $\bar{B}\in S_{loc}$ with $\inf_{B\in\mathcal{S}}\mathscr{L}_{loc}(B)\geq\mathscr{L}_{loc}(\bar{B})$ we get that $\bar{B}$ is a minimizer of $\mathscr{L}_{loc}$ in $\mathcal{S}_{loc}(A)$.
To understand the structure of minimizers, decompose $A,B$ using a reduced SVD to get $A=U_{A}\Sigma_{A}V_{A}^{\top},B=U_{B}\Sigma_{B}V_{B}^{\top}$. Notice that the constraint $AB^{+}B=A$ implies that $\mathrm{rank}(A)\leq\mathrm{rank}(B)$. Then the constraint is satisfied if and only if
$U_{A}\Sigma_{A}V_{A}^{\top}V_{B}\Sigma_{B}^{+}U_{B}^{\top}U_{B}\Sigma_{B}V_{B}^{\top}=U_{A}\Sigma_{A}V_{A}^{\top}V_{B}V_{B}^{\top}=U_{A}\Sigma_{A}V_{A}^{\top}.$
Multiplying on the left by $\Sigma_{A}^{+}U_{A}^{\top}$ we obtain
$V_{A}^{\top}V_{B}V_{B}^{\top}=V_{A}^{\top}$
and by multiplying on the right by $V_{A}$ we get
$\underbrace{(V_{A}^{\top}V_{B})}_{=:Q}(V_{B}^{\top}V_{A})=QQ^{\top}=I.$
Concerning the function value, by substituting $V_{A}^{\top}V_{B}=Q$ we see that
$\begin{split}\mathscr{L}_{loc}(U_{B}\Sigma_{B}V_{B})&=\frac{1}{2}\|U_{A}\Sigma_{A}V_{A}^{\top}V_{B}\Sigma_{B}^{+}U_{B}^{\top}\|^{2}=\frac{1}{2}\|U_{A}\Sigma_{A}Q\Sigma_{B}^{+}U_{B}^{\top}\|^{2}=\\ &=\frac{1}{2}\mathrm{tr}(U_{A}\Sigma_{A}Q\Sigma_{B}^{+}U_{B}^{\top}U_{B}(\Sigma_{B}^{+})^{\top}Q^{\top}\Sigma_{A}^{\top}U_{A}^{\top})=\\ &=\frac{1}{2}\mathrm{tr}(\Sigma_{A}Q\Sigma_{B}^{+}(\Sigma_{B}^{+})^{\top}Q^{\top}\Sigma_{A}^{\top})=\frac{1}{2}\|\Sigma_{A}Q\Sigma_{B}^{+}\|^{2}=\\ &=\frac{1}{2}\sum_{i=1}^{r}\sum_{j=1}^{m}s_{i}(A)^{2}s_{j}(B)^{-2}Q_{ij}^{2}=\frac{1}{2}\sum_{j=1}^{m}s_{j}(B)^{-2}\sum_{i=1}^{r}s_{i}(A)^{2}Q_{ij}^{2}.\end{split}$ (19)
Assume without loss of generality that the singular values of both $A,B$ are ordered in a descending manner. Now, since $QQ^{\top}=I$, this is a classical form of Brockett’s cost function [4], whose solution is given by $Q^{*}=[I_{r},0_{r,m-r}]$ by using Theorem 2.1 in [23], which is a generalization of the classical Ky-Fan trace inequality. Then, we have
$\min_{s_{j}(B)>0}\frac{1}{2}\sum_{j=1}^{r}\frac{s_{j}(A)^{2}}{s_{j}(B)^{2}},\quad\text{s.t.}\,s_{j}(B)^{2}\leq Cs_{j}(A)^{2},\forall j=1,\dots,r.$
which is minimized by $s_{k}(B)^{2}=Cs_{k}(A)^{2}$. The optimal function value is given by
$\mathscr{L}_{loc}(U_{B}\Sigma_{A}^{*}V_{A}^{\top})=\frac{1}{2}\sum_{j}\frac{s_{j}(A)^{2}}{Cs_{j}(A)^{2}}=\frac{1}{2C}\operatorname{rank}(A).$
Therefore, we have that minima are attained with $V_{B}(Q^{*})^{\top}=V_{A}$ for $Q^{*}=[I_{r},0_{r,m-r}]\in\mathbb{R}^{r\times m}$, $\Sigma_{B}\propto\mathrm{blockdiag}(\Sigma_{A},0_{m-r})$. This implies $V_{B}(Q^{*})^{\top}Q^{*}=V_{A}Q^{*}$. Therefore, the $r$ dominant right eigenvectors of $B$ coincide with those of $A$. Since the last mrm-r singular values of $B$ are zero, we can use the reduced SVD form and delete the last mrm-r columns of $V_{B}$, therefore obtaining $V_{B}=V_{A}$ and $\Sigma_{B}\propto\Sigma_{A}$. Moreover, $U_{B}$ is free, therefore all the optimal points are
$B^{*}\propto OU_{A}\Sigma_{A}V_{A}^{\top}=OA,\quad O\in\mathbb{R}^{m\times n},O^{\top}O=I\,(\text{since}\,n\leq m)$
and the optimal value is given by
$\mathscr{L}_{loc}(B^{*})=\frac{\operatorname{rank}(A)^{2}}{2C}.$
Now, to prove the existence of the path $\gamma(t)$, we assume $n=m=N$ and, without loss of generality, we consider $C=1$. From Equation˜19 it is sufficient to find a path for the singular values $(s_{i}(t))$, and for the matrix $Q(t)\in O_{N}(\mathbb{R})$ such that they make $\mathscr{L}_{loc}$ decrease and reach the optimal value.
The existence of a descent path in $Q$, for fixed $\Sigma_{A},\Sigma_{B}^{+}$, is guaranteed by Lemma˜B.2. Then, we can assume w.l.o.g. that $Q=I$ by assuming that the singular values of $A$ and $B$ are both ascendingly ordered. So we concatenate the path in $Q$ with the following one for the singular values of $B$:
$s_{i}(B)(t)=(1-t)s_{i}(B_{0})+ts_{i}(A).$ (20)
The path in $Q$ decreases $\mathscr{L}$ by Lemma˜B.2. Additionally, by taking the derivative in $t$ of $\mathscr{L}_{loc}$ along the path (20), from (19) we have
$\frac{d}{dt}\mathscr{L}_{loc}(OU_{A}\Sigma_{B}(t)Q(1)^{\top}V_{A}^{\top})$ $=\frac{d}{dt}\sum_{i,j=1}^{N}s_{i}^{2}(A)s_{j}^{-2}(B)(t)Q(1)^{2}_{ij}$
$=-2\sum_{i,j=1}^{N}s_{i}(A)^{2}s_{j}(B)(t)^{-3}\delta_{ij}(s_{j}(A)-s_{j}(B)(0))$
$=-2\sum_{i=1}^{N}s_{i}^{2}(A)s_{i}(B)(t)^{-3}(s_{i}(A)-s_{i}(B)(0))\leq 0,$
where in the second equality we have used that $Q(1)=I$ and the last inequality holds because $s_{i}(B_{0})\leq s_{i}(A)$. This concludes the proof. ∎
Lemma B.2 (Descent curve on $O_{N}(\mathbb{R})$).
Let $D_{1},D_{2}\in\mathbb{R}^{N\times N}$ be fixed diagonal positive definite matrices, each of them with distinct eigenvalues, and define
$\phi:O_{N}(\mathbb{R})\to\mathbb{R},\quad\phi(Q):=\|D_{1}QD_{2}\|^{2}=\mathrm{tr}(D_{1}^{2}QD_{2}^{2}Q^{\top}).$
Then, for almost every $Q_{0}\in O_{N}(\mathbb{R})$ there exists an analytic curve $Q:[0,1]\to O_{N}(\mathbb{R})$ such that
  • $Q(0)=Q_{0}$,
  • $Q(1)\in\underset{Q\in O_{N}(\mathbb{R})}{\arg\min}\,\,\phi(Q)$,
  • $\frac{d}{dt}\phi(Q(t))\leq 0$.
Proof.
First, we observe that the structure of all stationary and in particular, of global minima and saddle points is studied in [3, Theorem 4]. From this last result all the minima are global minima, and all saddle points are strict under the hypothesis that all the eigenvalues of both $D_{1}$ and $D_{2}$ are distinct. Let us now consider the Riemannian gradient flow
$\dot{Q}=-\mathrm{grad}\,\phi(Q)=-QD_{2}^{2}Q^{\top}D_{1}^{2}Q+D_{1}^{2}QD_{2}^{2},$
where $\mathrm{grad}$ is the gradient induced by the Frobenius inner product on $O_{N}(\mathbb{R})$. We now use [2, Theorem 6.3], which guarantees almost sure (with respect to the initial condition) convergence of the gradient flow to the local minima. Since we know that all local minima are global, we get the desired thesis.
Moreover, the solution curve $Q(t)$ is analytic because the vector field $\mathrm{grad}\,\phi(Q)$ is analytic. ∎
Using Lemma˜B.1, it is possible to prove the main result:
Theorem (Theorem˜4.2).
Let $\sigma_{2},\dots,\sigma_{L}$ be a set of entrywise homeomorphisms, $\sigma_{1}$ entrywise nonlinear with $\sigma_{l}(x)=0$ if and only if $x=0$. Assume $K=n_{L}\leq\dots\leq n_{1}=N$. Let $\mathcal{S}\subset\mathbb{R}^{n_{1}\times N}\times\dots\times\mathbb{R}^{n_{L}\times N}$ be as in Definition˜4.1. Then, the global minima of the optimization problem
$\min_{Z\in\mathcal{S}}\mathscr{L}(Z)$
satisfy $\operatorname{\mathrm{DNC1}}$ for all intermediate layers $l=1,\dots,L$. Moreover, any global minimizer $\bar{Z}$ satisfies
$\bar{Z}^{l}\propto O_{l}\sigma_{l+1}^{-1}(\bar{Z}^{l+1}),\quad O_{l}\in\mathbb{R}^{n_{l}\times n_{l+1}},O_{l}^{\top}O_{l}=I,\quad\forall\,l=1,\dots,L-1.$
Proof.
Without loss of generality, we assume $C=1$ in Definition˜4.1. Moreover, we can exclude the case $Z^{L}\neq Y$, since in that case $\mathscr{L}=+\infty$; therefore we assume $Z^{L}=Y$. To understand the existence and the structure of the minimizers, we can proceed iteratively. To this end, denote by $\mathcal{S}_{l}=\mathcal{S}_{l}(Z^{l-1},Z^{l+1})=\{Z^{l}\text{ that satisfy (I), (II) and (III)}\}$ for fixed, $Z^{l-1}$ and $Z^{l+1}$. Note in particular that
$S_{l}(Z^{l-1},Z^{l+1})\subseteq\mathcal{S}_{loc}(\sigma_{l+1}^{-1}(Z^{l+1})).$ (21)
As a consequence we study the infimum in $Z^{1}$ and observe the following set of inequalities
$\inf_{Z^{1}\in\mathcal{S}_{1}}\mathscr{L}(Z^{L},\dots,Z^{1})$ $\geq\min_{Z^{1}\in\mathcal{S}_{loc}(\sigma_{2}^{-1}(Z^{2}))}\mathscr{L}(Z^{L},\dots,Z^{1})$ (22)
$=\frac{1}{2}\sum_{l=2}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{\operatorname{rank}(\sigma_{2}^{-1}(Z^{2}))}{2}$
$\geq\frac{1}{2}\sum_{l=2}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{\operatorname{rank}(Y)}{2},$
where we have used first (21), second, Lemma˜B.1, which proves the existence of the minimum in $\mathcal{S}_{loc}(\sigma_{2}^{-1}(Z^{2}))$ and characterizes its value, and third, condition (III) in Definition˜4.1, which guarantees the last inequality when applied to $\operatorname{rank}(\sigma_{2}^{-1}(Z^{2}))$. Next, we focus on the lower bound in (22) and define the following function, depending only on $Z^{2},\dots Z^{L}$:
$\mathscr{L}_{1}(Z^{L},\dots,Z^{2}):=\frac{1}{2}\sum_{l=2}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{\operatorname{rank}(Y)}{2}.$
Thus, we proceed optimizing $Z^{2}$ in $\mathscr{L}_{1}$ as we did for $Z^{1}$ in $\mathscr{L}$ and obtain
$\inf_{\begin{subarray}{c}Z^{1}\in\mathcal{S}_{1},\\ Z^{2}\in\mathcal{S}_{2}\end{subarray}}\mathscr{L}(Z^{L},\dots,Z^{2},Z^{1})$ $\geq\inf_{Z^{2}\in\mathcal{S}_{2}}\mathscr{L}_{1}(Z^{L},\dots,Z^{2})$ (23)
$\geq\min_{Z^{2}\in\mathcal{S}_{loc}(\sigma_{3}^{-1}(Z^{3}))}\mathscr{L}_{1}(Z^{L},\dots,Z^{2})$
$=\frac{1}{2}\sum_{l=3}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{\operatorname{rank}(\sigma_{3}^{-1}(Z^{3}))}{2}+\frac{\operatorname{rank}(Y)}{2}$
$\geq\frac{1}{2}\sum_{l=3}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{2\operatorname{rank}(Y)}{2}.$
We continue iteratively by defining every time
$\mathscr{L}_{i}(Z^{L},\dots,Z^{i}):=\frac{1}{2}\sum_{l=i+1}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}+\frac{i\operatorname{rank}(Y)}{2}.$ (24)
Finally, we obtain
$\inf_{Z^{i}\in\mathcal{S}_{i}\;\forall i=1,\dots,L}\mathscr{L}(Z^{L},\dots,Z^{2},Z^{1})\geq\frac{L}{2}\operatorname{rank}(Y).$
Now we claim that there exists some $\bar{Z}\in\mathcal{S}$ such that $\mathscr{L}(\bar{Z})=(L/2)\operatorname{rank}(Y)$. Note that, as a consequence of the claim, we have that the minimum in $\mathcal{S}$ exists and that all the inequalities in (22) and (23) are equalities. In particular, because of Lemma˜B.1 any minimum $\bar{Z}$ has to satisfy
$\bar{Z}^{l}=O_{l}\sigma_{l+1}^{-1}(\bar{Z}^{l+1})\qquad\forall l=1,\dots,L-1,$ (25)
for some sequence of matrices $O_{l}\in\mathbb{R}^{n_{l}\times n_{l+1}}$ such that $O_{l}^{\top}O_{l}=I$. We conclude by observing that $Y=Z^{L}$ satisfies DNC1 and so, applying (25) backwards from $L$ to $1$, also all the other layers $\bar{Z}^{l}$, at the minima, satisfy DNC1.
We are left to prove the claim. To this end, recall that $y_{j}=e_{i_{j}}$ for any $j$ and $\sigma_{l}(0)=0$ for any $l$. So, $\sigma_{L}^{-1}(y_{j})=\alpha_{L}e_{i_{j}}$ for any $j$ where $\alpha_{L}$ is a constant different from zero and independent of $j$. In particular we can consider $O_{L-1}^{T}=[\mathrm{Id},0]$ and $\bar{Z}^{L-1}=O_{L-1}\sigma_{L}^{-1}(Y)=\alpha[Y,0]^{T}$ and iteratively do the same for any $l$. The obtained $\bar{Z}$ is then included in $\mathcal{S}$ because each layer $Z^{l}$ and its counterimage $\sigma_{l}^{-1}(Z^{l})$ are a rescaled version of $Y$, therefore conditions (I),(II),(III) are satisfied. Moreover, by its definition, $\|\sigma_{l+1}^{-1}(\bar{Z}^{l+1})\bar{Z}^{l,+}\|^{2}=\operatorname{rank}(\sigma_{l+1}^{-1}(\bar{Z}^{l+1}))=\operatorname{rank}(Y)$, yielding $\mathscr{L}(\bar{Z})=(L/2)\operatorname{rank}(Y)$ and concluding the proof of the claim. ∎
B.1 Reachability of DNC configuration for almost every initial condition
A natural question arising from the last section is: from which initial configurations can the global minima be reached? In this section, we will prove not only that the $\operatorname{\mathrm{DNC1}}$ minima are global, but also that they are reachable from any interpolating initial condition in a related setting thanks to the result in Theorem˜4.3.
We are now ready for the main result about the reachability of rank-collapsed configurations through energy-decreasing paths:
Theorem (Theorem˜4.3).
Assume the intermediate widths in $\mathcal{S}$ satisfy the condition $K\leq n_{L}=n_{L-1}=\dots=n_{1}=N$ and that the $\sigma_{l}$’s are analytic. Let $Z(0)\in\mathcal{S}$ be a random initial interpolating condition, i.e., such that $Z^{L}(0)=Y$. Then, almost surely there exists a continuous path $\gamma(t)$ with $\gamma(0)=Z(0)$, $\gamma(1)$ global minimum of $\mathscr{L}$, and such that $\mathscr{L}\circ\gamma$ is continuous Lebesgue-almost everywhere (except for a finite set of times) and for all continuity times $s\geq t$ we have
$\mathscr{L}(\gamma(t))\geq\mathscr{L}(\gamma(s)).$
Moreover, all discontinuities $t_{1},\dots,t_{m}$ are removable and satisfy
$\underset{t\to t_{i}^{-}}{\lim}\mathscr{L}(\gamma(t))=\underset{t\to t_{i}^{+}}{\lim}\mathscr{L}(\gamma(t))>\mathscr{L}(\gamma(t_{i})),\quad\forall i=1,\dots,m.$
Proof.
The idea is to construct a concatenation of curves layer by layer, where the existence of a curve for the local problem is given in Lemma˜B.1. Moreover, we notice immediately that according to Lemma˜B.1, this can be done for Lebesgue almost every $Z^{l}(0)\in\mathcal{S}$.
Define therefore $\gamma_{1}(t)$ as in Lemma˜B.1, to move $Z^{1}(0)$ to $Z^{2}(0), OO$$^{\top}O=I$. We recall the notation $\mathscr{L}(Z)=\sum_{l=1}^{L-1}\mathscr{L}_{l}$, where $\mathscr{L}_{l}:=\frac{1}{2}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}$. For the second layer, concatenate this path with $\gamma_{2}^{2}(t)$ a path connecting $Z^{2}(0)$ with an optimum $Z^{2,*}(Z^{3})$ and $\gamma_{2}^{1}(t)\in Z^{1,*}(\gamma_{2}^{2}(t))=\gamma_{2}^{2}(t)$. For this second path, thanks to Lemma˜B.3 and the hypothesis that $\operatorname{rank}(\sigma_{2}^{-1}(Z^{2}(0)))=N$, we have that for all $t\in[0,1]$ but a finite set, $\operatorname{rank}(\sigma_{2}^{-1}(\gamma_{2}^{2}(t)))=N$, and therefore the function $\mathscr{L}_{2}(\gamma_{2}^{2}(t))+\operatorname{rank}(\sigma_{2}^{-1}(\gamma_{2}^{2}(t)))^{2}$ is continuous up to a finite number of removable discontinuities. In the discontinuities $t_{i}$, as shown in Lemma˜B.3, $\operatorname{rank}(\sigma_{2}^{-1}(\gamma_{2}^{2}(t_{i})))<N$, the jump is down. In the other points, we have
$\frac{d}{dt}\mathscr{L}(Z^{L}(0),\dots,\gamma_{2}^{2}(t),\gamma^{1}_{2}(t))=\frac{d}{dt}\mathscr{L}_{2}(\gamma_{2}^{2}(t))+\frac{d}{dt}\operatorname{rank}(\sigma_{2}^{-1}(\gamma_{2}^{2}(t)))^{2}=\frac{d}{dt}\mathscr{L}_{2}(\gamma_{2}^{2}(t))\leq 0,$
where the last is because the rank of $\sigma_{2}^{-1}(\gamma_{2}^{2}(t))$ is constant along the path $\gamma_{2}^{2}(t)$. Again, inductively we can define a path $\gamma_{3}^{3}$ which makes $\mathscr{L}_{3}$ decrease and $\gamma_{3}^{2}(t)\in Z^{2,*}(\gamma_{3}^{3}(t)),\gamma_{3}^{1}(t)=Z^{1,*}(\gamma^{2}_{3}(t))=\gamma^{2}_{3}(t)$. Along this path, again, up to a finite number of times in which there is a jump down, we have
$\frac{d}{dt}\mathscr{L}(Z^{L}(0),\dots,\gamma_{3}^{3}(t),\gamma_{2}^{3}(t),\gamma^{1}_{3}(t))=\frac{d}{dt}\mathscr{L}_{3}(\gamma_{3}^{3}(t))\leq 0.$
At layer $l$, we generally have paths $\gamma_{l}^{l^{\prime}}(t)$ for which the required decrease condition is satisfied. By concatenating all of them we have the required path $\gamma$. At the last path, the rank collapses, at $t=1$. For $t\in[0,1]$ we have (again thanks to Lemma˜B.3 up to a finite number of points $\{t_{1},\dots,t_{k}\}$) that
$\frac{d}{dt}\mathscr{L}(Y,\gamma_{L-1}^{L-1}(t),\dots,\gamma_{L-1}^{1}(t))=$ $\frac{d}{dt}\mathscr{L}_{L-1}(\gamma_{L-1}^{L-1}(t))$ (26)
$+\sum_{l=1}^{L-2}\frac{d}{dt}\operatorname{rank}(\sigma_{l}^{-1}(\gamma_{L-1}^{l}(t)))\leq 0,\,\,\,\text{for\, a.e.}\,t\in[0,1).$
At $t=1$ instead, the rank collapses for all layers to $K$, therefore again decreasing the objective. ∎
Lemma B.3 (Finitely-many zeros of determinant on analytic matrix curves).
Consider an entrywise analytic function $g$ and a piecewise analytic curve $A:[0,1]\to\mathbb{R}^{N\times N}$ with $\operatorname{rank}g(A(0))=N$. Then there exists at most a finite number of times $\{t_{1},\dots,t_{k}\}$ in which such $\operatorname{rank}g(A(t))\leq N-1$. In particular, the function $t\mapsto\operatorname{rank}(g(A(t)))$ has just removable discontinuities, i.e., left and right limits exist and $\underset{t\to t_{i}^{+}}{\lim}\operatorname{rank}(g(A(t)))=\underset{t\to t_{i}^{-}}{\lim}\operatorname{rank}(g(A(t)))=N>\operatorname{rank}(g(A(t_{i})))$.
Proof.
We can prove it without loss of generality for a fully analytic curve $A(t)$, and then the same result holds on all subintervals in which $A$ is analytic (note that there are finitely many of them).
The function $t\mapsto\det(g(A(t)))$ is analytic and not constantly zero since $\det g(A(0))\neq 0$, and therefore, thanks to Lemma˜A.3, it has a finite number of zeros in $[0,1]$. Therefore, there can be just a finite number of points in which $\operatorname{rank}g(A(t))\leq N-1$. Now, we will prove that the discontinuities are removable. In particular, let $\varepsilon>0$. Since $\operatorname{rank}(g(A(t)))=N$ for all $t\in[0,1]\setminus\{t_{1},\dots,t_{k}\}$, there exists $\delta>0$ such that no $t_{j}\in(t_{i}-\delta,t_{i})$. In $(t_{i}-\delta,t_{i})$, $\operatorname{rank}(g(A(t))\equiv N$ and therefore we proved that the left limit exists and it is equal to $N$. The same can be done for the right limit. Therefore the concatenation of the paths $\gamma$ is continuous and $\mathscr{L}\circ\gamma$ has a finite number of discontinuities $E:=\{t_{i}\}_{i}$. ∎
As a corollary of Theorem˜4.3 we have the following:
Corollary B.4 (No loss barrier for any finite stepsize).
In the setting of Theorem˜4.3, for Lebesgue-almost every $Z(0)\in\mathcal{S}$ there exists a sequence of curves $\gamma_{k}:[0,1]\to\mathcal{S}$ uniformly converging to the continuous $\gamma$ of Theorem˜4.3 such that the following hold for all $k\geq 1$:
  • $\gamma_{k}$ is piecewise constant in a set of disjoint intervals intervals $\{I_{j,k}:=[T_{j,k},T_{j+1,k})\}_{j=0,\dots,k-1}$ whose union is $[0,1]$, and with $0<T_{j+1,k}-T_{j,k}<\frac{1}{k}$ for all $j=0,\dots,k-1$;
  • $\gamma_{k}(0)=Z(0)$;
  • $\gamma_{k}(1)\in\underset{Z\in\mathcal{S}}{\arg\min}\,\mathscr{L}(Z)$;
  • $\mathscr{L}$ is decreasing, i.e.,
    $\mathscr{L}(\gamma_{k}(0))\geq\mathscr{L}(\gamma_{k}(T_{1,k})))\geq\mathscr{L}(\gamma_{k}(T_{2,k})))\geq\dots\geq\mathscr{L}(\gamma_{k}(T_{k,k})))\geq\mathscr{L}(\gamma_{k}(1)).$
Proof.
Consider the curve $\gamma$ from Theorem˜4.3 together with its finite number of discontinuities $E:=\{t_{i}\}_{i}$ in which
$\underset{t\to t_{i}^{-}}{\lim}\mathscr{L}(\gamma(t))=\underset{t\to t_{i}^{+}}{\lim}\mathscr{L}(\gamma(t))>\mathscr{L}(\gamma(t_{i})).$
Consider now a sequence of sets $(\{T_{j,k}\}_{j=0,\dots,k-1})_{k}$, where $0<T_{j+1,k}-T_{j,k}\leq\frac{1}{k}$ for all $j$, $T_{0,k}=0,T_{k-1,k}=1$ for all $k$, and such that $T_{j,k}\notin E$ for all $j,k$. Define now the sequence of paths
$\gamma_{k}(s)=\gamma\Bigl{(}\sum_{j=0}^{k-1}T_{j,k}1_{[T_{j,k},T_{j+1,k})}(s)\Bigr{)}.$
Then, by construction, we have $\gamma_{k}:[0,1]\mathcal{\to}S$, $\gamma_{k}(0)=Z(0),\gamma_{k}(1)$ global minimizer. Moreover, since $\gamma_{k}(T_{j,k})=\gamma(T_{j,k})$ and $T_{j,k}\notin E$ we have that the sequence $(\mathscr{L}(\gamma_{k}(T_{j,k})))_{j}=(\mathscr{L}(\gamma(T_{j,k})))_{j}$ is decreasing. Finally, we show uniform convergence to the $\gamma$ constructed above, by using the fact that $\gamma$ is piecewise analytic and therefore Lipschitz:
$\sup_{s\in[0,1]}\|\gamma_{k}(s)-\gamma(s)\|$ $=\max_{j=0,\dots,k-1}\sup_{s\in[T_{j,k},T_{j+1,k}]}\|\gamma_{k}(s)-\gamma(s)\|$
$=\max_{j=0,\dots,k-1}\sup_{s\in[T_{j,k},T_{j+1,k}]}\|\gamma(T_{j,k})-\gamma(s)\|$
$\leq L\max_{j=0,\dots,k-1}|T_{j+1,k}-T_{j,k}|\leq\frac{L}{k}\to 0.$ $(\gamma\text{ Lipschitz})$
Therefore $\gamma_{k}$ converges to $\gamma$ uniformly. ∎
Remark B.5 (Intuition on Corollary˜B.4).
We highlight that an interpretation of Corollary˜B.4, is in terms of numerical schemes. In particular, since $\gamma_{k}:[0,1]\to\mathcal{S}$ is piecewise constant, we can think of it as an ordered sequence of $k$ points $\Gamma_{k}:=(\gamma_{k}(0),\gamma_{k}(T_{1,k}),\dots,\gamma_{k}(T_{k-1,k}),\gamma_{k}(1))$, where $\gamma_{k}(0)=Z(0)$ and $\gamma_{k}(1)\in\underset{Z\in\mathcal{S}}{\arg\min}\,\mathscr{L}$. We know from Corollary˜B.4 that $0<T_{j+1,k}-T_{j,k}:=\Delta_{j,k}<\frac{1}{k}$, and we can think of $\Delta_{j,k}$ as the stepsize of an optimization scheme. With this heuristic intuition in mind, the result proven in the corollary essentially states that for Lebesgue-almost every initialization $Z(0)\in\mathcal{S}$, there exists a numerical scheme with arbitrarily small maximal stepsize such that $\mathscr{L}$ decreases along it. Notice that, if there was any loss barrier between $Z(0)$ the the global minimizers of $\mathscr{L}$ on $\mathcal{S}$, then this would not be possible: in particular, if there was a barrier between initialization and minimizers, then there would exist a small enough stepsize such that $\mathscr{L}$ would need to increase along the sequence.
B.2 Stability of the $\operatorname{\mathrm{DNC1}}$ configurations, proof of Theorem˜4.4
Proof.
We can without loss of generality assume $Z^{L}=Y$, otherwise the condition is trivially satisfied. Therefore, we can assume $\mathcal{S}$ contains the interpolating condition $Z^{L}=Y$ as a constraint. Define and consider a sequence $Z_{\lambda}\to Z^{*}\in\mathcal{S}$ for $\lambda\to\infty$, $Z_{\lambda}\in\mathcal{S}$. Since $\mathcal{S}$ contains the interpolating condition, we have $\mathscr{L}_{\lambda}(Z^{*})<+\infty$. Define $M:=\lim\inf_{\lambda\to+\infty}\mathscr{L}_{\lambda}(Z_{\lambda})=\lim_{m\to+\infty}\mathscr{L}_{\lambda_{m}}(Z_{\lambda_{m}})$ for a converging subsequence $\lambda_{m}$ (we can exclude the case $M=+\infty$ cause the required inequality is trivially satisfied). Then, thanks to the lower semicontinuity of $\mathscr{L}_{\lambda}$ we get
$\mathscr{L}(Z^{*})\leq\mathcal{\liminf}_{m}\mathscr{L}(Z_{\lambda_{m}})=\liminf_{m\to\infty}\frac{1}{2}\sum_{l=1}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1}_{\lambda_{m}})Z^{l,+}_{\lambda_{m}}\|^{2}\leq\liminf_{m}\mathscr{L}_{\lambda_{m}}(Z_{\lambda_{m}})<+\infty.$
As a recovery sequence for the $\limsup$ condition we take $Z_{\lambda}\equiv Z^{*}$, in fact we have
$\limsup_{\lambda\to\infty}\mathscr{L}_{\lambda}(Z^{*})=\lim_{\lambda\to\infty}\mathscr{L}_{\lambda}(Z^{*})=\mathscr{L}(Z^{*}).$
Moreover, let $\lambda_{n}\to+\infty$ with $Z_{n}\in\arg\min_{Z\in\mathcal{S}}\mathscr{L}_{\lambda_{n}}(Z)$ and that it admits an accumulation point $Z^{*}\in\mathcal{S}$. Then there exists a convergent subsequence $Z_{n_{j}}\to Z^{*}$. For every $(\mathscr{Z}_{j})$ recovery sequence for a generic $\mathscr{Z}$ we have, for the definition of $\Gamma$-convergence and the fact that $Z_{n_{j}}$ is a minimizer of $\mathscr{L}_{\lambda_{n_{j}}}$,
$\mathscr{L}(Z^{*})\leq\lim\inf_{j}\mathscr{L}_{\lambda_{n_{j}}}(Z_{n_{j}})\leq\lim\inf_{j}\mathscr{L}_{\lambda_{n_{j}}}(\mathscr{Z}_{j})\leq\lim\sup_{j}\mathscr{L}_{\lambda_{n_{j}}}(\mathscr{Z}_{j})\leq\mathscr{L}(\mathscr{Z}),$
which implies $\mathscr{L}(Z^{*})\leq\mathscr{L}(\mathscr{Z})$ for every $\mathscr{Z}\in\mathcal{S}$, therefore $Z^{*}$ is a minimizer. ∎
Remark B.6.
Notice that of interpretation of Theorem˜4.4 is the implicit bias of fully connected networks in the regime $\|Y\|\ll\|X\|$, i.e. the regime in which $\lambda\to+\infty$.
A second interpretation can be formalized as follows (and proved in the exact same way): consider the representation cost
$\mathscr{L}^{[l_{1},l_{2}]}_{L}(Z):=\begin{cases}+\infty,\quad\text{if}\,\,Z^{L}\neq Y\\ \frac{1}{L}\sum_{l=l_{1}}^{l_{2}}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2},\quad\text{otherwise}\end{cases}$
and consider the difference (without loss of generality, for $Z^{L}=Y$)
$\mathscr{L}^{[0,L-1]}_{L}-\mathscr{L}^{[l^{*},L-1]}_{L}=\mathscr{L}^{[0,l^{*}-1]}_{L}=\frac{1}{L}\sum_{l=0}^{l^{*}-1}\|\sigma_{l+1}^{-1}(Z^{l+1})Z^{l,+}\|^{2}\underset{\Gamma}{\to}0,\quad L\to+\infty$
for any finite $l^{*}$, which $\Gamma$-converges to zero for $L\to\infty$. Therefore, in the regime of large depth, even in the absence of the piramidal architecture on the backbone, one expects the minimizers to satisfy exactly $\operatorname{\mathrm{DNC1}}$ from the first layer $l^{*}$ for which the set of representable features $Z^{l^{*}}$ contains the minimizer predicted in Theorem˜4.2 (or alternatively, as a special case, when the set of representable matrices $Z^{l^{*}}$ is the whole $\mathbb{R}^{n_{l^{*}}\times N}$). This observation formalizes the theory in more general cases, essentially stating that for deep enough architectures, the structure of the minimizers is close to the one predicted in Theorem˜4.2, where $Z^{1}$ is substituted with the minimal $l^{*}$ such that $Z^{l^{*}}$ is “representative enough”.
Appendix C Global minima of representation cost for residual networks
For residual networks, the representation of the layer is given by
$Z^{l+1}=\sigma_{l+1}(W_{l+1}Z^{l})+Z^{l}$
and therefore we have
$W_{l+1}=\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+},$
giving the representation cost
$\mathscr{L}^{res}(Z):=\begin{cases}+\infty,\quad\text{if}\,\,f(\mathit{\Theta};X)\neq Y\\ \frac{1}{2}\|\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}\|^{2}+\frac{1}{2}\sum_{l=1}^{L-2}\|\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}\|^{2},\quad\text{otherwise}.\end{cases}$
The proof of the main result relies on the following lemma:
Lemma C.1 (Optimal single layer for ResNet representation cost).
Let $\mathscr{L}_{loc}^{res}(B)=\frac{1}{2}\|g(A-B)B^{+}\|^{2}$, where $A\in\mathbb{R}^{n\times N}$ is a fixed matrix and $g:\mathbb{R}\to\mathbb{R}$ is entrywise nonlinear with $g(x)=0\iff x=0$. Then, for $n\leq m$, we have that
$\min_{B\in\mathbb{R}^{m\times N}}\mathscr{L}_{loc}^{res}(B)\quad\textnormal{s.t.}\quad g(A-B)B^{+}B=g(A-B)$
is attained at the point
$B^{*}=A.$
Proof.
Notice that $\mathscr{L}_{loc}^{res}(B)\geq 0$ and $\mathscr{L}_{loc}^{res}(B)=0$ if and only if $g(A-B)B^{+}=0$. Then we have because of the constraint that
$0=g(A-B)B^{+}B=g(A-B).$
So the only point attaining the minimum satisfies g(AB)=0g(A-B)=0 which implies $B^{*}=A$. ∎
In this setting we have the following result about global minima:
Theorem C.2 ($\operatorname{\mathrm{DNC1}}$ is optimal for constrained representation cost).
Let $\sigma_{l}$ be a sequence of entrywise nonlinearities for which $\sigma_{l}(x)=0\iff x=0$ and assume intermediate widths satisfy the condition $K\leq n_{L}=n_{L-1}=\dots=n_{1}=N$. Then, the global minima of the optimization problem
$\min_{Z^{1},\dots,Z^{L}}\mathscr{L}^{res}(Z)$
$\textnormal{s.t.}\quad\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l}),\quad\forall\,l=1,\dots,L-2$
$\qquad\,\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}Z^{L-1}=\sigma_{L}^{-1}(Z^{L})$
$\qquad\,s_{i}(Z^{L-1})\leq Cs_{i}(\sigma_{L}^{-1}(Y)),\quad\forall i=1\,\dots,r_{L}$
satisfy $\operatorname{\mathrm{DNC1}}$ for all intermediate layers $l=1,\dots,L$.
Proof.
Let’s define with $\mathcal{S}$ the full constraint set and with $\mathcal{S}_{l}=\mathcal{S}_{l}(Z^{1},\dots,Z^{l-1},Z^{l+1},\dots,Z^{L})$ the constraint set on $Z^{l}$ with the other variables fixed. Notice that we can first minimize in $Z^{1}$ the problem
$\arg\min_{Z^{1}\in\mathcal{S}_{1}}\mathscr{L}^{res}(Z^{L},\dots,Z^{1})=\arg\min_{Z^{1}\in\mathcal{S}_{1}}\|\sigma_{2}^{-1}(Z^{2}-Z^{1})Z^{1,+}\|^{2}$
$\qquad\;\text{s.t.}\quad\sigma_{2}^{-1}(Z^{2}-Z^{1})Z^{1,+}Z^{1}=\sigma_{2}^{-1}(Z^{2}-Z^{1})$
whose minima is given by Lemma˜C.1 at $Z^{1,*}(Z^{2})=Z^{2}$ and minimal value equal to zero. Therefore, we can now minimize in $Z^{2}$ the representation cost
$\min_{Z^{2},\dots,Z^{L}}\mathscr{L}^{res}(Z^{L},\dots,Z^{2},Z^{1,*}(Z^{2}))=\frac{1}{2}\|\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}\|^{2}+\frac{1}{2}\sum_{l=2}^{L-1}\|\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}\|^{2}$
$\text{s.t.}\quad\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}Z^{l}=\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})\quad\forall\,l=2,\dots,L-1$
$\qquad\,\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}Z^{L-1}=\sigma_{L}^{-1}(Z^{L})$
$\qquad\,s_{i}(Z^{L-1})\leq Cs_{i}(\sigma_{L}^{-1}(Y)),\quad\forall i=1\,\dots,K.$
Notice that the situation now is the same as the initial one, but with one layer less. Therefore, we can do the same thing for $Z^{2}$ using Lemma˜C.1 (with $g=\sigma_{l}^{-1}$) to obtain $Z^{2,*}(Z^{3})=Z^{3}$. By iterating this procedure for all layers, we arrive at
$\min_{Z^{L-1},Z^{L}}\frac{1}{2}\|\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}\|^{2}$
$\text{s.t.}\quad Z^{L}=Y$
$\qquad\,\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}Z^{L-1}=\sigma_{L}^{-1}(Z^{L})$
$\qquad\,s_{i}(Z^{L-1})\leq Cs_{i}(\sigma_{L}^{-1}(Y)),\quad\forall i=1\,\dots,K.$
By using Lemma˜B.1 we get that
$Z^{L-1,*}=O_{l}\sigma_{L}^{-1}(Y),\quad O_{l}^{\top}O_{l}=I.$
Therefore, $Z^{L-1,*}$ satisfies $\operatorname{\mathrm{DNC1}}$ cause $\tilde{Y}=\sigma_{L}(Y)$ satisfies $\operatorname{\mathrm{DNC1}}$. Recursively backward, we have by Lemma˜C.1 that
$Z^{L-2,*}=Z^{L-1,*},$
which is therefore also collapsed. In the same way, by backsubstituting the optimal point, we get
$Z^{L-1,*}\quad\text{collapsed}$
$Z^{L-2,*}=Z^{L-1,*}$
$\vdots$
$Z^{1,*}=Z^{2,*}$
and therefore
$Z^{L-1,*}=Z^{L-2,*}=\dots=Z^{1,*}$
and so they all satisfy $\operatorname{\mathrm{DNC1}}$. ∎
Remark C.3.
(Stability of minima for ResNets) We highlight that the same $\Gamma$-convergence result as the one presented in Theorem˜4.4 can be stated also for ResNets, and the proof is exactly the same as the one presented in Section˜B.2. More precisely, we define
$\mathscr{L}_{\lambda}^{res}(Z)=\begin{cases}+\infty,\quad\text{if}\,\,\ Z^{L}\neq Y,\\ \frac{1}{2}\|\sigma_{L}^{-1}(Z^{L})Z^{L-1,+}\|^{2}+\frac{1}{2}\sum_{l=1}^{L-2}\|\sigma_{l+1}^{-1}(Z^{l+1}-Z^{l})Z^{l,+}\|^{2}+\frac{1}{\lambda}\|\sigma_{1}^{-1}(Z^{1})X^{+}\|^{2},\quad\text{otherwise}.\end{cases}$ (27)
Then for $\lambda\to+\infty$, $\mathscr{L}_{\lambda}^{res}\to\mathscr{L}^{res}$ in the $\Gamma$-convergence sense, analogously to Theorem˜4.4. This will also be showcased numerically in Appendix˜D.
Appendix D Additional numerical experiments
In this section, we include additional numerical results. All experiments were performed on a Single NVIDIA A100 (80GB) GPU.
D.1 Additional values of $\delta$ for the singular value convergence
$\delta=10$ $\delta=8$
Refer to caption Refer to caption
$\delta=5$ $\delta=2$
Refer to caption Refer to caption
$\delta=1$ $\delta=10^{-1}$
Refer to caption Refer to caption
$\delta=10^{-2}$ $\delta=0$
Refer to caption Refer to caption
Figure 4: Singular values of trained networks against the predicted one from Lemmas˜B.1 and 4.3 for different values of $\delta=\frac{1}{\lambda}$ in Equation˜12. As we can observe, the singular values are converging to the predicted ones for $\delta\to 0$ as shown in Theorem˜$4.$
D.2 Residual networks
In this section we will show numerical results concerning Theorem˜C.2. The setting is the same of Section˜5, with $L=6$ layers (first and last feedforward, and the $4$ in the middle are residual). From Figure˜5, we can observe the evolution of the norm of the weights during optimization, also for varying $\delta=\frac{1}{\lambda}$ in Equation˜12. As we can see, for $\delta$ going to zero, all the weight matrices converge to zero as expected from Theorem˜C.2, but the first and the last ones, which are feedforward layers.
$\delta=8$ $\delta=5$
Refer to caption Refer to caption
$\delta=2$ $\delta=1$
Refer to caption Refer to caption
$\delta=10^{-1}$ $\delta=10^{-2}$
Refer to caption Refer to caption
$\delta=10^{-3}$ $\delta=0$
Refer to caption Refer to caption
Figure 5: Norm of the intermediate weights matrices for a for $L=6$ layer ResNet during training. As discussed in Remark˜C.3, we observe convergence to the structured minima discussed in Theorem˜C.2 for $\delta=\frac{1}{\lambda}$ (as in Equation˜12).
D.3 Visual representation of neural collapse
In this section we will include the evolution of the gram matrices $G^{l}=Z^{l,\top}Z^{l}$ for all layers at different phases of training. Notice that $\operatorname{\mathrm{DNC1}}$ happens if and only if $G^{l}$ has block structure. In fact, we have the following characterization:
Proposition D.1 (Characterization of $\operatorname{\mathrm{DNC1}}$ features through covariance).
The following are equivalent:
  1. 1.
    $Z\in\mathbb{R}^{n\times N}$ satisfied NC1;
  2. 2.
    $G=Z^{\top}Z=\left[\begin{array}[]{c|c|c|c|c}\alpha_{11}1_{N_{1}}1^{\top}_{N_{1}}&\alpha_{12}1_{N_{1}}1^{\top}_{N_{2}}&\alpha_{13}1_{N_{1}}1^{\top}_{N_{3}}&\dots&\alpha_{1K}1_{N_{1}}1^{\top}_{N_{K}}\\ \hline\cr\alpha_{21}1_{N_{2}}1^{\top}_{N_{1}}&\alpha_{22}1_{N_{2}}1^{\top}_{N_{2}}&\alpha_{23}11^{\top}&\dots&\alpha_{2K}11^{\top}\\ \hline\cr\vdots&\dots&\ddots&\dots&\alpha_{K-1,K}11^{\top}\\ \hline\cr\alpha_{K1}1_{N_{K}}1^{\top}_{N_{1}}&\alpha_{K,2}11^{\top}&\alpha_{K,3}11^{\top}&\dots&\alpha_{K,K}11^{\top}\end{array}\right]$ has block structure.
Proof.
Assume $Z$ satisfies NC1. For simplicity, we will denote with $z_{i}^{k}$ the $i$-th vector of the $k$-th class. Then we have
$Z^{\top}Z=\left[\begin{array}[]{c}1_{N_{1}}z^{1,\top}\\ 1_{N_{2}}z^{2,\top}\\ \vdots\\ 1_{N_{K}}z^{K,\top}\\ \end{array}\right]\left[\begin{array}[]{c}z^{1}1_{N_{1}}^{\top},z^{2}1_{N_{2}}^{\top},\dots z^{K}1_{N_{K}}^{\top}\end{array}\right]=\Bigl{(}\langle z^{i},z^{j}\rangle 1_{N_{i}}{1}_{N_{j}}^{\top}\Bigr{)}_{i,j=1,\dots,K},$
which has the required structure of $Z^{\top}Z$ with $\alpha_{ij}=\langle z^{i},z^{j}\rangle$. Vice versa, assuming that $Z^{\top}Z$ has the required structure, we have
$\alpha_{kk}=\langle z_{i}^{k},z_{j}^{k}\rangle,\quad\forall i,j=1,\dots,N_{k},\forall k=1,\dots,K.$
This implies that
$0=\langle z_{i}^{k},z_{j}^{k}-z_{m}^{k}\rangle,\quad\forall i,j,m=1,\dots,N_{k}.$
This, in turn, implies by definition of orthogonal projection that
$z_{i}^{k}=P^{\perp}_{\mathbb{A}^{k}}(0),\quad\forall i=1,\dots,N_{k},\forall k=1,\dots,K,$
where $\mathbb{A}_{k}=\text{Aff}(\{z_{j}^{k}\}_{j})$ is the affine subspace passing trhough the points of class $k$. By uniqueness of the orthogonal projection we get
$z_{i}^{k}=z_{j}^{k}=P_{\mathbb{A}^{k}}^{\perp}(0),\quad\forall i,j=1,\dots,N_{k},\forall k=1,\dots,K,$
which means $Z$ satisfies NC1. ∎
Given this premise, in the following experiments we will plot the Gram matrices $G_{l}=Z^{l,\top}Z^{l}$ for any $l=0,\dots,L-1$ in the case $L=15$ layer neural network trained in the setting of Section˜5. The emergence of the block structure is depicted in Figure˜6.
Moreover, we sampled a random plane in each feature space, i.e., for any $l$ we sample a random orthogonal matrix $B^{l}\in\mathbb{R}^{2\times n_{l}},B^{l}B^{l,\top}=I_{2}$. Figure˜7 shows the projection of the feature matrices $Z^{l}$ on the plane spanned by the rows of $B^{l}$, i.e. $B^{l}Z^{l}$, for which we plot all the columns as points in $\mathbb{R}^{2}$, allowing us to pictorially see the deep-neural (rank) collapse in all intermediate layers in this random subspace.
Epoch 0 Epoch 100
Refer to caption Refer to caption
Epoch 200 Epoch 400
Refer to caption Refer to caption
Epoch 700 Epoch 1200
Refer to caption Refer to caption
Figure 6: Emergence of the block structure in $Z^{l,\top}Z^{l}$ characterizing $\operatorname{\mathrm{DNC1}}$ during training. Layers for each epoch are ordered from left to right and from top to down.
Epoch 0 Epoch 100
Refer to caption Refer to caption
Epoch 200 Epoch 300
Refer to caption Refer to caption
Epoch 700 Epoch 1200
Refer to caption Refer to caption
Figure 7: Neural collapse of features in a random $d=2$ dimensional subspace, where each color represents a class of MNIST. Layers for each epoch are ordered from left to right and from top to down.